### v1

# Adv-BMT: Bidirectional Motion Transformer for Safety-Critical Traffic Scenario Generation

Abstract
Scenario-based testing is essential for validating the performance of autonomous driving (AD) systems. However, such testing is limited by the scarcity of long-tailed, safety-critical scenarios in existing datasets collected in the real world. To tackle the data issue, we propose the Adv-BMT framework, which augments real-world scenarios with diverse and realistic adversarial interactions. The core component of Adv-BMT is a bidirectional motion transformer (BMT) model to perform inverse traffic motion predictions, which takes agent information in the last time step of the scenario as input, and reconstruct the traffic in the inverse of chronological order until the initial time step. The Adv-BMT framework is a two-staged pipeline: it first conducts adversarial initializations and then inverse motion predictions. Different from previous work, we do not need any collision data for pretraining, and are able to generate realistic and diverse collision interactions. Our experimental results validate the quality of generated collision scenarios by Adv-BMT: training in our augmented dataset would reduce episode collision rates by 20% compared to previous work. The code will be made available at https://metadriverse.github.io/adv-bmt/.
1 Introduction
In recent years, autonomous driving (AD) agents have achieved unprecedented performance in urban environments dosovitskiy2017carlaopenurbandriving ; li2022metadrivecomposingdiversedriving ; wu2024metaurbanembodiedaisimulation . However, handling corner traffic situations, especially collision scenarios, remains a major challenge. One cause is that safety-critical scenarios are missing from real-world driving datasets due to high costs and risks of data collections. Without enough collision training data, it is hard for the autonomous driving (AD) planners and prediction models to learn safe driving in challenging and risky scenarios. This motivates the need for simulating collision traffic that emulates real-world accidents. According to rempe2022generatingusefulaccidentpronedriving , useful safety-critical scenarios should be realistic and diverse, as well as challenging. Previous works rempe2022generatingusefulaccidentpronedriving ; zhang2023catclosedloopadversarialtraining ; stoler2025sealsafeautonomousdriving ; ransiek2024goosegoalconditionedreinforcementlearning leveraged learned real-world traffic priors to constraint adversarial trajectories, and optimized predictions on collision-encouraging objectives. However, from our observations and evaluations, generated behaviors from these methods have a lack of diversity and a lower collision success rates (attacking efficiency).
To improve upon previous work, we propose the Adv-BMT framework for realistic and diverse collision augmentations from real-world driving data. Adv-BMT samples various collision situations from the input driving log, and then reconstruct complete adversarial trajectories accordingly. This design enhances the diversity of predicted collision interactions, since now it augments different collision behaviors for each driving log. Besides, there is no need to include collision data in trainings. We further implement a rule-based rejection sampling mechanism to filter out candidate trajectories from unsatisfactory adversarial initializations, and maintain realistic collision interactions. In the core of Adv-BMT is the bidirectional motion transformer (BMT) model. Similar to existing multi-agent simulators and prediction models wu2024smartscalablemultiagentrealtime ; seff2023motionlmmultiagentmotionforecasting ; zhou2024behaviorgptsmartagentsimulation , BMT tokenizes continuous agent trajectories into discrete actions, and performs next-token predictions in an auto-regressive manner. Different from existing models, BMT learns to predict both future and history motion for all agents, conditioning on one-step current states. This is achieved by processing two sets of motion tokens from the input scenario, with one set of tokens used for forward prediction (future motion) and one for reverse prediction (history motion).
As a summary, Adv-BMT is a two-staged pipeline: first, it initializes diverse collision frames between an new adversary agent (ADV) and ego vehicle; then, it reconstructs the adversarial trajectories via BMT’s reverse predictions. In the output, ADV is added to the original scenario, maintaining realistic interactions with surrounding traffic. An overview of our framework is illustrated in Fig. 1. We summarize our contributions as follows: (1) We introduce the bidirectional motion transformer with temporally reversible motion tokenizations; (2) We develop Adv-BMT for realistic and diverse safety-critical traffic simulations; (3) We leverage Adv-BMT in a closed-loop setting to dynamically create challenging environments for reinforcement learning agents.
2 Related Work
Driving Motion Predictions and Simulations
Recent work have utilized auto-regressive sequence modeling on discretized motion representations via transformer-based models. In previous works, Trajeglish philion2024trajeglishtrafficmodelingnexttoken discretizes motion representations using a k-disk based tokenization scheme to represent position and angle differences for relative movements; but the interaction performance is sensitive to agent ordering, and agents decoded first might not be as reactive to the rest of the predicted agents. MTR++ shi2024mtrmultiagentmotionprediction directly represents motion on continuous space in Gaussian mixture distributions. The design is efficient for multi-modal generations but prediction diversities across modes rely on the learned intention queries. MotionLM seff2023motionlmmultiagentmotionforecasting and SMART wu2024smartscalablemultiagentrealtime model motion actions that represents trajectory deltas on decoder-only transformers. BehaviorGPT zhou2024behaviorgptsmartagentsimulation performs next-patch predictions with future motion chunks, instead of single-step predictions, and achieves higher inference efficiency and reduces compounding errors.
Safety-Critical Scenario Generations
Previous works leverage learned traffic prior to generate realistic safety-critical interactions. STRIVE rempe2022generatingusefulaccidentpronedriving enhances diversity through modeling traffic a latent space using variational autoencoders but requires computationally expensive per-scenario optimizations. CAT zhang2023catclosedloopadversarialtraining utilizes the augmentation framework into MetaDrive li2022metadrivecomposingdiversedriving simulator as a closed-loop reinforcement learning environment. Upon CAT, SEAL stoler2025sealsafeautonomousdriving uses a skill-based adversarial policy for human-like driving guided by collision-related objectives. MixSim Suo_2023_CVPR generates realistic reactions by exploring future goals and re-simulates the inferred trajectories using a trained policy. GOOSE ransiek2024goosegoalconditionedreinforcementlearning uses a goal-conditioned reinforcement learning approach, which parameterizes trajectories and goal constraints to generate scenario-level, safety-critical interactions. AdvSim wang2023advsimgeneratingsafetycriticalscenarios directly perturbs actor trajectories using a kinematic model and optimizes via a black-box adversarial loss. Recent work ding2020multimodalsafetycriticalscenariosgeneration uses a conditional normalizing flow to model the distribution of real-world safety-critical trajectories. SafeSim chang2024safesimsafetycriticalclosedlooptraffic proposes a diffusion model with a test-time collision-sensitive guidance loss to control the collision type and adversarial agent selections; however its generation quality relies on initial collections of collision data. CrashAgent li2025crashagentcrashscenariogeneration and LCTGen tan2023languageconditionedtrafficgeneration leverage free-form texts as inputs and extract embeddings to parameterize scene initializations and agent driving directions.
Figure 1: Overview of Adv-BMT framework. From the input scenario, Adv-BMT first samples diverse collision states between ego agent and a new adversarial agent, then reconstruct complete adversarial trajectories via BMT’s reverse motion predictions. A rule-based reject sampling mechanism is used for filtering unsatisfactory collision initializations.
3 Methodology
In this section, we formulate forward and reverse motion predictions. We then introduce bidirectional motion transformer (BMT) model. Last but not the least, we demonstrate details of Adv-BMT.
3.1 Problem Formulation
Consider a traffic scenario with maximumly $N$ agents and time horizon $T$. The trajectory of agent $i$ is denoted by $\boldsymbol{\tau}^{i}=\{{\tau}^{i}_{0},{\tau}^{i}_{1},\dots,{\tau}^{i}_{T}\}$, where each state $\tau^{i}_{t}\in\mathbb{R}^{d}$ includes position, velocity, and heading at time $t$. We define a prediction direction indicator $D\in\{\text{Forward},\text{Reverse}\}$, for future or past motion prediction within $T$ time steps. For each scenario we construct a sequence of motion tokens:
$\mathbf{Z}^{i}=\{z^{i}_{1},\dots,z^{i}_{T}\},\quad z^{i}_{k}=\begin{cases}\phi(\tau^{i}_{k},\tau^{i}_{k-1}),&D=\text{Forward}\\ \phi(\tau^{i}_{T-k},\tau^{i}_{T-k+1}),&D=\text{Reverse}\end{cases}$ (1)
where $\phi(\cdot)$ denotes the motion tokenization function which we discuss in section 3.2. The goal is to model the distribution of agent motion tokens at the $k$-th step, conditioned on the predicted motion token sequence, current state, and scene context $\mathcal{M}$:
$P(z^{1}_{k},z^{2}_{k},\dots,z^{N}_{k}\mid\mathbf{z}^{1}_{1:k-1},\mathbf{z}^{2}_{1:k-1},\dots,\mathbf{z}^{N}_{1:k-1},\mathcal{M},D)$ (2)
where $\mathbf{Z}^{i}_{0:k-1}$ denotes history token sequence for agent $i$. The decoder autoregressively selects the next action token $z^{i}_{k+1}$ by attending to past tokens and scene token embedding $\mathcal{M}$. Each selected action in forward prediction maps to a position in future time-step, similar to a standard motion predictor; selected token sequence in reverse prediction can be reconstructed as history trajectories. Together, BMT is able to perform both future and past motion predictions as shown in Fig. 4.
Figure 2: BMT model architecture.
3.2 Bi-direction Motion Tokenizations
Token space
BMT’s bidirectional motion tokens are derived from a simplified bicycle dynamics model. Both forward and reverse tokens are defined over the same shared token space—a set of discrete bins of acceleration and yaw rate pairs. We uniformly quantize the control space of accelerations $a\in[-a_{\max},a_{\max}]$ and yaw rates $\delta\in[-\delta_{\max},\delta_{\max}]$ into $K$ bins each, yielding a total of $K^{2}$ discrete motion tokens, where $a_{\max}=10m/s$, $\delta_{\max}=\frac{\pi}{2}$, and $K=33$.
Control token simulation
Given a motion control token, BMT simulates it and reconstructs to a new agent state using midpoint integration:
$\text{Recons}(z,\tau,\Delta t)=\left(x+\bar{v}\cdot\cos(\bar{\theta})\cdot\Delta t,\;y+\bar{v}\cdot\sin(\bar{\theta})\cdot\Delta t,\;\theta+\omega\cdot\Delta t,\;v+a\cdot\Delta t\right)$ (3)
where $\tau=(x,y,\theta,v)$ denotes the current agent state with position $(x,y)$, heading $\theta$, and speed $v$; $z=(a,\omega)$ is the motion token containing acceleration $a$ and yaw rate $\omega$; $\Delta t$ is the time interval; $\bar{v}=v+\frac{1}{2}\cdot a\cdot\Delta t$ is the midpoint speed; and $\bar{\theta}=\theta+\frac{1}{2}\cdot\omega\cdot\Delta t$ is the midpoint heading.
Ground-truth tokens
For obtaining ground truth motion tokens during trainings, we simulate all $K^{2}$ candidate motion tokens starting from each agent state at time $t$, then select the token whose predicted next state $\hat{\tau}_{t+1}$ best matches the ground truth $\tau_{t+1}$. This process maps the continuous trajectory into a sequence of discrete motion tokens $\{z_{0},z_{1},\dots,z_{T}\}$ for each predicted agent. Forward and Reverse predictions share the same motion token space but differ in token interpretations. In reverse predictions, we invert the temporal order of the trajectory and treat the final state (e.g., collision initializations from Adv-BMT) as the initial; at each time step, we simulate each motion token (by inverting the time step, $\Delta t\to-\Delta t$), and find the token that best reconstructs the previous state $\tau_{t}$ from $\tau_{t+1}$. This generates a motion token sequence in reverse time direction $\{\tilde{z}_{0},\tilde{z}_{1},\dots,\tilde{z}_{T}\}$, where $\tilde{z}_{0}$ corresponds to the final state in forward time.
3.3 Token Matching with minimized Contour Error
For either reverse or forward predictions, we build motion tokens step-by-step in the order of prediction time, starting from the first prediction step and reconstructing till the last prediction step. With a 0.5-second time interval, a search is conducted from all candidate token vocabularies, and selects the best candidate that minimizes the distance between the ground-truth and reconstructed contours for the $i$-th candidate action for agent $n$:
$a_{n}^{*}=\arg\min_{i}\left(\mathcal{E}_{n}^{(i)}=\frac{1}{P}\sum_{p=1}^{P}\left\|\mathbf{c}_{n}^{(i,p)}-\hat{\mathbf{c}}_{n}^{(p)}\right\|_{2}\right)\quad\text{where }0\leq i<K^{2}$ (4)
where $\mathbf{c}_{n}^{(i,p)}\in\mathbb{R}^{2}$ and $\hat{\mathbf{c}}_{n}^{(p)}\in\mathbb{R}^{2}$ represent the $p$-th polygon corner of the reconstructed and ground-truth contours, each with totally $P$ contour points; $a_{n}^{*}$ is the index of the selected candidate token.
3.4 Bidirectional Motion Transformer (BMT)
The BMT architecture and training process is shown in Fig. 2, along with ground-truth token matching for forward predictions or reverse predictions.
Architecture
BMT has a scene encoding component used to obtain embeddings for scenario contexts with separate embeddings for map polylines, traffic lights, and agent initial states. Then, we use Fourier-encoded edge features tancik2020fourierfeaturesletnetworks to represent the spatial and directional information between these encoded entities, which are then passed to the transformer encoder with self-attention layers for the relational embeddings.
The prediction decoder predicts subsequent motion tokens in an autoregressive manner, with only initial agent information for the first frame, along with the scene embedding obtained from the Scene Encoder. The motion decoder incorporates self-attention over the initial agent token embeddings, and three relation computations separately: agent-to-agent (a2a), agent-to-time (a2t), and agent-to-scene (a2s), with each relation embedding then passed to its cross-attention layers. The output agent embeddings are concatenated and repeated a number of times. The output agent motion embeddings are mapped to the vocabulary of discretized motion tokens through MLPs.
Training
In training, BMT predicts the sequence of motion tokens and use the cross-entropy objective to match the joint action distribution of the observed behaviors in the dataset during training:
$\max_{\pi_{\theta}}\;\mathbb{E}_{\mathcal{D}}\left[\sum_{t=1}^{T}\sum_{i}^{N}\log\pi_{\theta}\left(z^{i}_{t}\mid\mathbf{z}^{i}_{1:k-1},\mathcal{M},D\right)\right]$ (5)
During inference, BMT generate motion tokens autoregressively, and uses nucleus sampling (top-$p$ sampling) to enhance diversity and maintain plausibility. To prevent the exposure bias during auto-regressive inference, BMT rolls out upon chosen motion tokens instead of the ground-truth token sequence.
3.5 Adv-BMT
The overview of Adv-BMT is illustrated in Fig.1. The framework mainly consists of three steps: (1) it adds an adversarial agent (ADV) with a specified collision state into the current scenario; (2) it predicts reversely for the adversarial trajectory; (3) it performs rule-based checks and reject physically implausible ones.
Diverse Adversarial Initializations
While existing works select a convenient neighbor agent and modify the behavior to attack the ego agent, Adv-BMT inserts new agents (ADV) with diverse collision initializations for different opponent interactions. We visualize an example result in 4. Formally, we define the collision state as to include the position, time, velocity, and heading at the collision step. The collision time is sampled from the first second to the last timestep of the ego trajectory length. Similarly, ADV’s collision headings are randomly sampled. ADV’s collision position can be calculated from collision heading and the ego vehicle’s collision position, with the ADV contour intersecting with the ego vehicle’s contour. Last but not least, collision velocity is also calculated from sampling a variable offset from ego vehicle’s speed at collision step.
Reverse Predictions
We use the sampled collision states as inputs of BMT for reverse predictions, and obtain ADV’s trajectory from predictions to augment the original scene with a new adv agent. During BMT inference, we teacher-force the ground-truth token sequence of ego agent and traffic agents, to keep the originality of the scenario as much as possible (since BMT could predict agents to driving on a different lane choice, which makes it a totally different scenario). By teacher-forcing traffic vehicle agents, BMT reconstructs the adversarial agent trajectory to fit in the input scenario. Reconstructing entirely new scenarios is out of the scope of our current work, but BMT is able to realize both normal and safety-critical generations through forward and reverse predictions.
Rule-based Rejection Samplings
We design Adv-BMT to have diverse initializations, which do not guarantee the realism of the collision outcomes. To address this issue, we implement a rule-based rejection mechanism for ADV candidates. We first measure the driving distance and average speed of ADV candidates; if it moves too short and mostly wanders at the designated position waiting for the ego vehicle, then it is considered invalid. Meanwhile, we check the max curvature (the rate of change for heading): given a candidate ADV trajectory, we compute the curvature constraint using:$\kappa_{t}={\Delta\theta_{t}}/{\Delta s_{t}},$ where $\Delta\theta_{t}$ is the absolute heading change between time steps, and $\Delta s_{t}$ is the displacement between consecutive positions. Adv-BMT rejects a prediction if $\max_{t}(\kappa_{t})>\kappa_{\text{threshold}},$where $\kappa_{\text{threshold}}=0.8$ is a predefined curvature limit that we found useful. We check the curvature to ensure it does not exceed our defined threshold, and reject candidate ADVs with unrealistic changes in headings. With our straightforward rule-based rejection sampling mechanism, we are able to maintain only realistic collision events between the ADV and the ego vehicle.
Figure 3: BMT predicts both future and history agent trajectories.
Figure 4: Adv-BMT generates diverse collision scenarios.
Forward refinement for Reactive Traffic
In real-world collision events, neighbor vehicles would be affected by the collision vehicles and would have reactions like sudden lane changes and brakings. With this concern, we implement Bi-directional Adv-BMT, with a forward refinement to re-simulate traffic agents upon predicted ADV trajectory. In forward refinement, all traffic agents would be predicted and modified in the output scenario, with teacher-forcing motions of collision participants (ego and ADV agents). In experiments, we refer to it as Adv-BMT (Bi) to indicate bi-directions.
4 Experiments
In section 4.1, we evaluate BMT model on motion prediction performance. In section 4.2, we compare Adv-BMT with baseline methods. In section 4.3 we train a reinforcement learning agent in different safety-critical augmented datasets. For augmentations, we use ground-truth traffic data from the Waymo Open Motion Datasets (WOMD) ettinger2021largescaleinteractivemotion .
4.1 Realism of Adv-BMT
Settings
We measure BMT on four different settings: Forward prediction with or without agent history, Reverse prediction without agent history, and bi-directional predictions. We measure on open-loop scenario generation metrics with 6 prediction modes on each scenario: scenario final displacement errors (SFDE) which measures the average displacement error at the final step on the best prediction mode, and scenario average displacement errors (SADE) which measures the averages according to all time steps on the best mode. We also measure Final Displacement Diversity (FDD) and average displacement diversity (ADD) for prediction diversities across different prediction modes: FDD measures the spread of final positions and ADD measures trajectory variations over time. Finally, we measure the overall collision rates for realistic interactions. The results are presented in Table 1.
Table 1: Evaluation for BMT on scenario realism on different prediction modes.
Context SFDE $\downarrow$ SADE $\downarrow$ minSFDE $\downarrow$ minSADE $\downarrow$ FDD $\uparrow$ ADD $\uparrow$ Collision
Forward 1s 3.043 1.019 1.908 0.693 9.241 3.240 0.026
Forward 0s 3.180 1.297 2.144 0.927 10.263 4.177 0.028
Reverse 0s 3.411 1.380 2.158 0.975 8.234 2.913 0.031
Bidirections 0s 3.059 2.111 1.831 1.536 7.618 6.270 0.035
Analysis
Results from Table 1 indicate that BMT performs realistic scenario generations in both forward predictions and reverse predictions. Furthermore, by conditioning on a 1-second history, the forward prediction achieves better SADE and SFDE scores, indicating better temporal consistencies. Due to design of Adv-BMT with agent initializations at the last step, it is impossible for reverse predictions to access agent histories. Compared to single-pass prediction, bi-directional prediction exhibit lower prediction accuracy in ADE metrics, indicating a harm in the realism of generations.
In terms of prediction diversities, Forward prediction exhibits higher FDD and ADD scores than reverse prediction. This indicates that predicting histories is more constrained, but predicting future motions encourages more explorations for variant directions. This further validates the necessity of Adv-BMT’s diverse collision initialization design to force different collision outcomes. Despite these differences, the overall collision rate remains reasonable and comparable for all prediction modes, and indicates that both prediction modes effectively generate realistic interactions.
Figure 5: Comparison among adversarial behaviors of different methods.
4.2 Safety-critical Evaluations
In following experiments, We compare Adv-BMT with three baseline methods for safety-critical augmentations, namely STRIVE rempe2022generatingusefulaccidentpronedriving , CAT zhang2023catclosedloopadversarialtraining , and SEAL stoler2025sealsafeautonomousdriving . We evaluate the quality and efficiency of generated collisions among all methods. We measure the adversarial attack success rate (SDC-ADV collision rate), collision rates between adversary agents and traffic agents (ADV-BV Collision rate), indicating interaction realism. Additionally, we evaluate the difference between the predicted distribution and the ground-truth distribution on velocities (VEL JSD) and accelerations (ACC JSD). These two metrics also indicate the realism of generated adversarial agent motions.
Settings
We randomly select 500 WOMD scenarios, each as a base environment for adversarial augmentations. For each scenario, each method predicts six different adversarial trajectories. To be consistent with baseline methods in this evaluation only, we force Adv-BMT to choose an existing neighbor agent and modify its behavior as the predicted adversary.
Table 2: Evaluation for Safety-critical Scenarios
[…]
CAT zhang2023catclosedloopadversarialtraining 0.470 0.271 0.137 0.480 0.0 0.0
SEAL stoler2025sealsafeautonomousdriving 0.582 0.261 0.239 0.561 0.0 0.0
STRIVE rempe2022generatingusefulaccidentpronedriving 0.534 0.358 0.117 0.432 0.004 0.001
Adv-BMT 1.000 0.391 0.142 0.435 0 2.344
Adv-BMT (Bi) 0.997 0.373 0.188 0.534 0.038 2.257
Quantitative Results
Table 2 shows the quantitative results for scenario generation metrics of all methods. Two settings of Adv-BMT outperform baselines in terms of adversarial attack success rate. This validates Adv-BMT’s design of collision initialization + reverse predictions, which couldn’t be achieved by other methods. While Adv-BMT generates highly adversarial behaviors, it also preserves diversity of traffic interactions compared to baselines. Note tha FDD and SDD metrics suggest that our baselines generate nearly the same adversarial trajectory on the given scene, while Adv-BMT is able to generate diverse behaviors at different prediction mode. The JSD metrics suggests slightly better realism on velocities and accelerations for baselines compared to Adv-BMT, showing the trade-off between diversity and velocity/acceleration realism. BMT model is able to achieve realistic motion predictions indicted by Waymo Open Sim Agent Challenge metrics montali2023waymoopensimagents , which we append in our appendix. Overall, our safety-critical evaluations indicate that Adv-BMT is an efficient framework for realistic, diverse, and safety-critical generations.
Visualizations
We simulate Adv-BMT scenarios in MetaDrive li2022metadrivecomposingdiversedriving simulators using both BEV and 3D rendering. In Fig. 4, we show diverse Adv-BMT adversarial behaviors in the collision directions on the same ego agent in an intersection scenario from WOMD. Adversarial agents (in red) come from different lanes and collide with the ego agent (in green) at different angles and facets. Besides, the generated adversarial agent follows the traffic rules and maintains realistic driving patterns. Apart from diverse collision directions, Adv-BMT generates collisions at diverse timings. Adv-BMT adversarial agents cover collisions with the ego before entering the intersection, during the middle of the intersection, and after turning to the main lane. With a limited number of real-world maps and logs, Adv-BMT is able to generate diverse collision behaviors and timings, making it suitable for AD testing and adversarial training. A visual comparison among Adv-BMT and baseline methods is shown in Fig. 5, where other methods fail to generate an adversarial trajectory but Adv-BMT is able to insert a realistic ADV and triggers collisions .
Generation Speeds
Adv-BMT and CAT generate safety-critical scenarios significantly faster than other baselines. Adv-BMT generates one scenario in 1.02s compared to SEAL’s 2.36s and STRIVE’s 9.53s, achieving great efficiency gains. With details shown in Table 7 in Appendix.
4.3 Adversarial Learning
To validate the value of Adv-BMT scenarios in downstream autonomous driving (AD) tasks, we train a reinforcement learning agent within augmented scenarios with collision interactions generated by Adv-BMT and baseline methods. To determine the quality of augmented training scenarios, we measure both driving performance and safety performance of the learned AD agent compared to learning in original training set. In our experiment, we conduct two sets of trainings. (1) open-loop training: agent is trained in the fixed training set with adversarial scenarios generated based on ground-truth ego trajectories. (2) closed-loop training: an adaptive adversarial agent attacks on current ego agent based on its recent rollout trajectory records. The adaptive adversarial agent’s motion is generated by Adv-BMT and baseline. Results of both trainings are shown in table 3 and table 4 respectively.
Settings
The training set contains 500 real-world scenarios randomly selected from WOMD training set. Each waymo scenario is augmented with one collision scenario, so that the ratio between waymo and safety-critical scenes is 1:1. For each method, we augments the original dataset into a new training set. For the adaptive adversarial learning experiment, we implement an adaptive generator for Adv-BMT and CAT zhang2023catclosedloopadversarialtraining (discard other methods due to lower inference time). We train a Twin Delayed DDPG (TD3) agent for 1 million steps using 8 random seeds to ensure robustness. We conduct the training in the MetaDrive simulator and use the TD3 algorithm for reinforcement learning (hyper-parameters listed in Appendix). We measure the average reward, average cost, average route completion rate (Compl.), and average episode cost (cost sum) for driving performance measure. We further conducted ablation studies with results shown in Appendix.
Quantitative Results
To evaluate the impact of adversarial training using Adv-BMT-generated scenarios, we assess policy performance across two distinct validation environments: (1) 100 Waymo Validation Environments, which consist of unmodified real-world driving scenarios from WOMD validation set, and (2) 100 Adv-BMT Environments, which is the augmented collision scenarios from the 100 validation scenes. Table 3 presents the evaluation results, demonstrating how training with Adv-BMT improves policy resilience in both standard and adversarial settings. A detailed analysis of these results follows in the subsequent sections.
Table 3: Open-loop learning evaluation.
(a) Evaluation in the Waymo Validation Environments
Training Scenarios Reward $\uparrow$ Cost $\downarrow$ Compl. $\uparrow$ Coll. $\downarrow$ Cost Sum $\downarrow$
Waymo ettinger2021largescaleinteractivemotion 32.03 $\pm$ 4.27 0.39 $\pm$ 0.07 0.72 $\pm$ 0.05 0.14 $\pm$ 0.02 1.41 $\pm$ 0.35
CAT zhang2023catclosedloopadversarialtraining 30.37 $\pm$ 3.89 0.39 $\pm$ 0.05 0.71 $\pm$ 0.05 0.14 $\pm$ 0.02 1.73 $\pm$ 0.39
STRIVE rempe2022generatingusefulaccidentpronedriving 31.30 $\pm$ 3.59 0.40 $\pm$ 0.04 0.73 $\pm$ 0.05 0.13 $\pm$ 0.03 1.51 $\pm$ 0.40
SEAL stoler2025sealsafeautonomousdriving 29.94 $\pm$ 5.14 0.39 $\pm$ 0.05 0.71 $\pm$ 0.04 0.12 $\pm$ 0.02 1.63 $\pm$ 0.44
Adv-BMT 31.47 $\pm$ 3.21 0.38 $\pm$ 0.03 0.73 $\pm$ 0.04 0.11 $\pm$ 0.02 1.35 $\pm$ 0.40
Adv-BMT (Bi) 33.22 $\pm$ 1.83 0.36 $\pm$ 0.03 0.74 $\pm$ 0.03 0.12 $\pm$ 0.02 1.39 $\pm$ 0.22
(b) Evaluation in the Adv-BMT Environments
Training Scenarios Reward $\uparrow$ Cost $\downarrow$ Compl. $\uparrow$ Coll. $\downarrow$ Cost Sum $\downarrow$
Waymo ettinger2021largescaleinteractivemotion 37.01 $\pm$ 6.16 0.64 $\pm$ 0.09 0.60 $\pm$ 0.07 0.30 $\pm$ 0.02 2.96 $\pm$ 0.63
CAT zhang2023catclosedloopadversarialtraining 36.77 $\pm$ 4.95 0.62 $\pm$ 0.05 0.62 $\pm$ 0.05 0.29 $\pm$ 0.02 3.09 $\pm$ 0.56
STRIVE rempe2022generatingusefulaccidentpronedriving 37.72 $\pm$ 5.38 0.63 $\pm$ 0.06 0.63 $\pm$ 0.06 0.29 $\pm$ 0.04 2.92 $\pm$ 0.68
SEAL stoler2025sealsafeautonomousdriving 35.74 $\pm$ 6.36 0.67 $\pm$ 0.08 0.60 $\pm$ 0.06 0.31 $\pm$ 0.01 2.97 $\pm$ 0.34
Adv-BMT 37.33 $\pm$ 3.57 0.62 $\pm$ 0.03 0.63 $\pm$ 0.04 0.25 $\pm$ 0.05 2.41 $\pm$ 0.43
Adv-BMT (Bi) 39.55 $\pm$ 2.94 0.59 $\pm$ 0.04 0.65 $\pm$ 0.02 0.27 $\pm$ 0.04 2.74 $\pm$ 0.54
Table 4: Closed-loop learning results
(a) Waymo Validation Environments
Adaptive Generator Reward $\uparrow$ Cost $\downarrow$ Comp. $\uparrow$ Coll. $\downarrow$ Cost Sum $\downarrow$
CAT 32.15 $\pm$ 2.89 0.38 $\pm$ 0.03 0.74 $\pm$ 0.04 0.10 $\pm$ 0.00 2.02 $\pm$ 0.24
Adv-BMT 33.13 $\pm$ 4.11 0.39 $\pm$ 0.03 0.74 $\pm$ 0.03 0.09 $\pm$ 0.00 1.25 $\pm$ 0.52
(b) Adv-BMT Environments
Adaptive Generator Reward $\uparrow$ Cost $\downarrow$ Comp. $\uparrow$ Coll. $\downarrow$ Cost Sum $\downarrow$
CAT 39.47 $\pm$ 3.88 0.62 $\pm$ 0.05 0.63 $\pm$ 0.04 0.22 $\pm$ 0.02 2.51 $\pm$ 0.45
Adv-BMT 40.40 $\pm$ 6.39 0.57 $\pm$ 0.04 0.63 $\pm$ 0.05 0.22 $\pm$ 0.04 2.48 $\pm$ 0.97
Analysis
1) Improved Policy Robustness in Adversarial Settings. Table 3 presents the evaluation results comparing policies trained with Adv-BMT-generated scenarios against baselines in both real-world and safety-critical validation environments. Adv-BMT demonstrates superior performance in reducing collision rates while maintaining high completion rates. In the collision validation environments, Adv-BMT-trained policies achieve the lowest collision rate while preserving competitive completion rates. This highlights the effectiveness of Adv-BMT in exposing policies to diverse and high-risk scenarios, leading to improved safety performance. Furthermore, Adv-BMT achieves the lowest sum cost, suggesting that policies trained on Adv-BMT-generated data are more efficient in handling safety-critical interactions. Besides, the results from adaptive generators shows an slight improvement in performance compared to training in fixed scenario sets.
In the open-loop learning, among the baselines, STRIVE achieves the highest reward (37.72) in collision environments, but this comes at the cost of a higher collision rate (0.29). SEAL and CAT exhibit higher costs and collision rates, which suggests that adversarial training leads to suboptimal behaviors. In contrast, agent training with Adv-BMT attacks results in improvement on both safety and task completion. We conduct an ablation study on Adv-BMT plus the forward refinement method, and results indicate a slightly better performance in average rewards, and average cost compared to Adv-BMT.
From closed-loop learning shown in table 4, we can see that agent learned with Adv-BMT environments has a significantly lower driving trajectory costs and higher rewards compared to baseline methods. With last step costs, collision rates, and driving completion rate remain consistent with the baseline. Comparing across tables with open-loop evaluations, we can see improvements over all metrics including rewards, completion rates, collision rates, and trajectory costs. This validates the value of diverse safety-critical scenarios which is provided in the closed-loop setting.
5 Conclusion and Limitations
Conclusion
This work introduces Adv-BMT, a novel safety-critical scenario generation framework that leverages BMT model to bring diverse and realistic adversarial collisions into real-world traffic scenarios. Adv-BMT consists of three key components: (1) collision initialization, which samples a potential collision outcome as the last frame; (2) reverse prediction, which generates the adversarial pre-collision trajectories; (3) rule-based filtering, which ensures adversarial agents’ motions remain realistic and feasible. We evaluate Adv-BMT across multiple benchmarks and experiments and demonstrate that training with generated adversarial scenarios improves AD agent performance in both held-out log-replay and collision test sets with a clear margin.
Limitations
Following limiations of our current work will be potentially addressed in the future: (1) we only consider adversarial behaviors among vehicle agents, but not among other agent types like pedestrains and bicycles. (2) we only evaluate our adversarial scenarios in RL policies, but adversarial scenarios can also be leveraged in human-in-the-loop learning.
Appendix A Model Details
A.1 BMT Architecture
Scenario Format and Preprocessing
The model takes real-world scenarios with the format from ScenarioNet li2023scenarionetopensourceplatformlargescale , which unifies real-world datasets including caesar2020nuscenesmultimodaldatasetautonomous  caesar2022nuplanclosedloopmlbasedplanning  wilson2023argoverse2generationdatasets .
The model performs a series of preprocessing on both static map features and dynamic agent trajectories. We compute the global boundary to extract a consistent map center and heading. Each map feature is decomposed into a sequence of vectorized segments, where each vector is represented by a start and end point in 3D coordinates. These vectors are then augmented with directional relative positions, segment headings, and lengths. Semantic attributes are binary indicators that encode feature types. The map feature types (lane, crosswalk, broken line, yellow line, stop sign, etc.) are encoded for semantic information. Features are then centralized to map center. Traffic light states are also encoded and aligned with the steps. Similarly, agent trajectories are also centralized for agent feature encode, with information of (position, heading, velocity, shape, and type) extracted across time for temporal sequences. Besides, there is reordering for ego agent and agents to be predicted. We extract a 16-dimensional state vector at each timestep. The preprocessing ensures all trajectories are represented in a consistent spatial frame and translated to an ego-centric coordinate system.
Tokenization
We formulate motion prediction as an auto-regressive next-token prediction problem, where each motion token corresponds to a discretized control input for a fixed time interval. Tokenization process maps continuous agent motion (acceleration and yaw rate) into discrete 2D bins. The tokenization process considers candidate tokens sampled from a fixed grid of bins, and simulates the resulting motion over a short duration, and then selects the best-matching action by minimizing the contour alignment error between the predicted agent shape and the ground-truth position and heading. We adopt a simplified version of the bicycle model to parameterize agent motion using longitudinal acceleration and yaw rate within predefined bounds: acceleration is limited to the range of $[-10,10]$ m/s2, and the yaw rate is constrained to $[-\pi/2,\pi/2]$ rad/s. With predicted motion token sequences, the trajectory can be reconstructed by mapping the token back to the acceleration and yaw angle change.
In both forward and reverse directions, motion tokens are decoded into continuous trajectories using the same tokenizer. Given an initial state $\tau_{t}=(x_{t},y_{t},\theta_{t},v_{t})$, the forward decoding process simulates the agent’s next state $\tau_{t+1}$ by applying a tokenized control action $z_{t}=(a_{t},\delta_{t})\in\mathcal{A}$, where $\mathcal{A}$ denotes the discrete token space. This is repeated autoregressively over the sequence of predicted tokens to reconstruct the full trajectory $\boldsymbol{\tau}_{t:t+T}$. In contrast, reverse decoding starts from a known future state $\tau_{t+1}$, the model evaluates all possible token candidates $z_{t}\in\mathcal{A}$, simulates their inverse dynamics using $\Delta t\to-\Delta t$, and selects the token that best reconstructs the preceding state $\tau_{t}$. The ability to operate in both directions is a key distinction of our approach: forward prediction enables open-loop simulation of future behaviors, while reverse prediction allows us to trace back from a desired outcome (e.g., a collision state) to plausible initiating actions.
Decoding Architecture.
The decoder follows a GPT-like structure composed of stacked cross-attention layers. Each layer integrates three structured attention modules: agent-to-agent (A2A), agent-to-temporal (A2T), and agent-to-scene (A2S). These modules attend over dynamically constructed graphs defined by spatial or temporal adjacency. Relational information across modalities is captured via multiple embeddings: Agent-to-Agent Relation Embedding, Agent-to-Time Relation Embedding, and Agent-to-Scene Relation Embedding, each encoding context-specific spatial information. For token construction, several embeddings are used, including the Agent Type Embedding, Agent Shape Embedding, Agent ID Embedding, and the Motion Token Embedding which maps discrete control tokens. A Continuous Motion Feature Embedding is applied to embed acceleration and yaw rate attributes. Auxiliary embeddings include the Special Token Embedding for indicating sequence boundaries, and the Backward Prediction Indicator Embedding to distinguish between forward and backward prediction modes. For each attention edge, a relative relation embedding is computed using a Fourier encoder and added to the key and/or value vectors. The input agent token $\mathbf{X}\in\mathbb{R}^{B\times T\times N\times d}$ (batch size $B$, time steps $T$, number of agents $N$, hidden dimension $d$) is progressively updated across layers by aggregating contextual features from other agents, their temporal history, and relevant map elements.
Motion Token Embedding
Specifically, each input agent token to the BMT motion decoder is constructed by summing of embeddings of:
  • a.
    the motion token from the previous step (representing discretized acceleration and yaw rate),
  • b.
    agent shape (length, width, height),
  • c.
    agent type (e.g., vehicle, pedestrian),
  • d.
    agent identifier (embedded optionally),
  • e.
    special token type (e.g., <start>, <end>, or padding),
as well as a continuous motion delta feature embedded via a Fourier encoder. These components are projected into the same hidden dimension and summed to form the input motion token embedding.
Prediction Heads and Outputs
After processing through all decoding layers, the final hidden state for each valid token is passed through a two-layer MLP head to produce logits over the motion token space:
$\texttt{MLP}(\mathbf{h})=W_{2}\cdot\phi(W_{1}\cdot\mathbf{h})\in\mathbb{R}^{K^{2}},$
where $\phi(\cdot)$ denotes the GELU activation and $K^{2}$ is the number of discrete motion tokens (from a $K\times K$ acceleration–yaw bin grid).
The resulting logits are used to predict the next motion token at each time step. During inference, we generate motion tokens using nucleus (top-$p$) sampling.
Trajectory Reconstruction
Our model makes a prediction in the interval of 5 time steps (0.5 seconds). To simulate the effect of a motion token over a fixed time step $\Delta t=0.5\,\mathrm{s}$, we adopt midpoint integration based on a simplified bicycle model. In forward predictions, given a current state $\mathbf{s}_{t}=(x_{t},y_{t},\theta_{t},v_{t})$, the model computes the next speed and heading as $v_{t+1}=v_{t}+a\cdot\Delta t$ and $\theta_{t+1}=\theta_{t}+\omega\cdot\Delta t$. The average speed and heading are then defined as $\bar{v}=\frac{v_{t}+v_{t+1}}{2}$ and $\bar{\theta}=\left(\frac{\theta_{t}+\theta_{t+1}}{2}\right)$. In reverse prediction, the process is reverted. In the backward direction, the process is inverted. Given a future state $\mathbf{s}_{t+1}$, the model enumerates all possible token candidates and inverts the dynamics: $v_{t}=v_{t+1}-a\cdot\Delta t$ and $\theta_{t}=\theta_{t+1}-\omega\cdot\Delta t$. The average quantities $\bar{v}$ and $\bar{\theta}$ are computed similarly and used to derive the previous position:
$x_{t}=x_{t+1}-\bar{v}\cdot\cos(\bar{\theta})\cdot\Delta t,\quad y_{t}=y_{t+1}-\bar{v}\cdot\sin(\bar{\theta})\cdot\Delta t.$
Decoder Training Loss
The decoder produces a logit tensor $\hat{\mathbf{z}}\in\mathbb{R}^{B\times T\times N\times|\mathcal{A}|}$, where $|\mathcal{A}|$ is the number of motion tokens (i.e., discretized acceleration–yaw pairs). The supervision target is the ground-truth token sequence $\mathbf{z}^{*}\in\mathbb{N}^{B\times T\times N}$, derived by tokenizing agent trajectories. A binary mask $\mathbf{m}\in\{0,1\}^{B\times T\times N}$ specifies which tokens are valid and should contribute to the training loss. The training objective is computed over all valid entries using the cross-entropy loss:
$\mathcal{L}_{\text{main}}=\frac{1}{\sum_{b,t,n}m_{b,t,n}}\sum_{b,t,n}m_{b,t,n}\cdot\text{CE}(\hat{z}_{b,t,n},z^{*}_{b,t,n}),$
where CE denotes the standard cross-entropy loss between the predicted logits and the ground-truth discrete token.
Reverse Prediction
During reverse prediction, the model measures metrics separately for forward and reverse token predictions. Let $\mathbf{b}\in\{0,1\}^{B\times T\times N}$ be a binary indicator for whether each token comes from reverse prediction. Then we compute separate metrics:
$\text{Accuracy}^{\text{reverse}}$ $=\frac{\sum m_{b,t,n}\cdot b_{b,t,n}\cdot\mathbf{1}[\hat{z}_{b,t,n}=z^{*}_{b,t,n}]}{\sum m_{b,t,n}\cdot b_{b,t,n}},$
$\text{Entropy}^{\text{reverse}}$ $=\frac{1}{\sum m\cdot b}\sum m_{b,t,n}\cdot b_{b,t,n}\cdot\mathcal{H}(\hat{z}_{b,t,n}),$
with analogous expressions for forward prediction (i.e., for $1-b_{b,t,n}$).
Metrics
To measure the quality and diversity of the model’s predictions during training, we track the perplexity:
$\text{Perplexity}=\exp\left(-\sum_{a\in\mathcal{A}}\bar{p}_{a}\log(\bar{p}_{a}+\epsilon)\right),\quad\text{where}\quad\bar{p}_{a}=\frac{1}{M}\sum_{i=1}^{M}\mathbf{1}[\hat{z}_{i}=a],$
and $M$ is the number of valid tokens. We also track the number of distinct tokens used by both predictions and ground truth:
$\text{Cluster}=\sum_{a\in\mathcal{A}}\mathbf{1}[\bar{p}_{a}>0].$
Total Loss
The total loss is the sum of all enabled components:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{main}}+\lambda_{\text{map}}\mathcal{L}_{\text{map}}+\lambda_{\text{tg}}\mathcal{L}_{\text{tg-total}},$
with default weights $\lambda_{\text{map}}=\lambda_{\text{tg}}=1$.
Table 5: BMT Model Parameters.
Component Parameters Size (MB)
Scene Encoder 902,080 3.44
   Map Polyline Encoder 22,656 0.09
   Traffic Light Embedding MLP 1,024 0.00
   Scene Relation Embedding 117,184 0.45
   Scene Transformer Encoder 744,448 2.84
   Scene Encoder Output Projection 16,512 0.06
   Scene Output Pre-Normalization 256 0.00
Motion Decoder 4,385,025 16.73
   Multi-Cross Attention Decoder 2,881,536 10.99
   Motion Prediction Head 157,121 0.60
   Motion Prediction Pre-Normalization 256 0.00
   Agent-to-Agent Relation Embedding 418,432 1.60
   Agent-to-Time Relation Embedding 418,432 1.60
   Agent-to-Scene Relation Embedding 117,184 0.45
   Agent Type Embedding 640 0.00
   Motion Token Embedding 139,520 0.53
   Agent Shape Embedding 17,152 0.07
   Agent ID Embedding 16,384 0.06
   Continuous Motion Feature Embedding 217,600 0.83
   Special Token Embedding 512 0.00
   Reverse Prediction Indicator Embedding 256 0.00
Total 5,287,105 20.17
A.2 Training Details
Our model has 5.2 million trainable parameters, with detailed break downs indicatd in Table 5. We trained BMT model on the training set of Waymo Open Motion Datasets ettinger2021largescaleinteractivemotion . WOMD contains 480K real-world traffic with each scenario of length 9 seconds; traffic are composed by agents of vehicle, pedestrian, and bicycle; Each scenario comes with a high-fidelity road map. During training, we use 8 NVIDIA RTX A6000 GPUs for our model training and fine-tunings. We trained BMT in two stages, each with hyper-parameters indicated in Table 6. In the first stage, we pre-trained BMT for forward prediction only with 1 million steps. Then, we fine-tuned BMT with reverse motion prediction in fine-tuning with totally 1.5 million steps. We use AdamW optimizer for learning rate scheduling.
Table 6: BMT Training settigs.
Forward Prediction
Hyper-parameter Value
Training steps 10E6
Batch sizes 2
Training Time (h) 185
Sampling Topp 0.95
Sampling temperature 1.0
Learning Rates 3E-4
Reverse Prediction
Hyper-parameter Value
Training steps 15E6
Batch size 2
Training Time (h) 310
Sampling Topp 0.95
Sampling temperature 1.0
Learning Rates 3E-4
Figure 6: Diverse adversarial behaviors generated by Adv-BMT.
Appendix B Experiment Details
Training Environment
We conduct our reinforcement learning experiments using the MetaDrive ScenarioEnv li2022metadrivecomposingdiversedriving , which provides standardized driving environments for training and evaluating autonomous agents. Each environment encodes sensor observations including LiDAR-based surroundings and physical dynamics. Specifically, the observation space consists of three key components: (i) Ego state, which contains the ego vehicle’s current physical state such as speed, heading, and steering; and (ii) surroundings, which encodes nearby traffic objects.
Actions are continuous and correspond to low-level vehicle control commands. The agent outputs a 2D normalized vector, which is then mapped to steering angle, throttle (acceleration), and brake signal. The environment includes a compositional reward structure combining driving progress, collision penalties, and road boundary violations. Driving reward is measured by forward lane progress, while penalties are applied for collisions with other vehicles or drifting off-road.
Hyper-parameters
The settings of our open-loop and closed-loop adversarial RL experiments are shown in table 7. Note that in our closed-loop learning, Adv-BMT takes one frame of agent information as input for adversarial generations, whereas all baseline methods take one second agent history.
Table 7: RL training settings.
Adv-BMT
Hyper-parameter Value
Scenario Horizon 9s
History Horizon 0s
Collision Step 1s–9s
Prediction Mode 8
Policy Training Steps 10E6
TD3
Hyper-parameter Value
Discounted Factor 0.99
Train Batch Size 1024
Learning Rate 1E-4
Policy Delay 200
Target Network 0.005
Appendix C Quantitative Results
C.1 Waymo Open Sim Agents Challenge
Table 8: WOSAC Evaluation results of BMT.
Metrics Reverse Forward
Linear speed 0.375 0.393
Linear acceleration 0.394 0.405
Angular speed 0.441 0.428
Angular acceleration 0.594 0.593
Distance to nearest object 0.405 0.388
Collision 0.521 0.951
Time to Collide 0.840 0.840
Distance to road edge 0.675 0.683
Offroad 0.564 0.934
Realism 0.554 0.753
Kinematic 0.451 0.455
Interactive 0.566 0.801
Map 0.596 0.862
minADE 2.148 1.344
Metametric 0.554 0.753
We evaluated BMT resutls on 400 WOMD validation scenarios using Waymo Open Sim Agents Challenge (WOSAC) 2024 montali2023waymoopensimagents . Evaluation results are summarized in Table 8. Within the metrics, for minADE metrics, smaller values indicate better more accurate perdictions, and for the rest metrics, the larger score means better performance. Within the results, we can see that forward prediction achieves much better performance compared to reverse prediction across all metrics, except similar performance in Angular speed, angular acceleration, distance to nearest object, and TTC. In these mentioned metrics. Note that the training time for forward prediction and reverse prediction are similar (10E6 and 15E6 steps respectively). The WOSAC results indicate that our BMT model is better at predicting future motion than predicting history motion. We will submit to the official WOSAC learderboard for further reference in the future.
Appendix D Qualitative Results
Visualizations
In Fig. 6, we demonstrate six pairs of qualitative results by Adv-BMT. In different scenarios, our opponent behavior exihibits in different category of safety-critical driving patterns. This means that our augmented opponent agents have plausible interactions with realistic driving trajectories. Besides, in Fig. 7, we demonstrate diverse collisions from just one single scenario. This is the power of Adv-BMT compares to other baseline work, which generates the same or similar adversarial behavior with the same driving log.
Figure 7: Diverse adversarial behaviors generated on one driving log by Adv-BMT.
Demo Video
We submit a video within our supplementary materials. Here we put visualizations with case studies through animated visualizations of Adv-BMT scenarios, which simulates different types of vehicles, pedestrians and bicycle agents.
Appendix E Broader Impacts
Our work introduces a novel model for generating safety-critical traffic scenarios, aiming to improve the safety reliability and driving robustness of AD systems. By modeling both forward and reverse motion trajectories, our framework enables controllable and diverse simulation of rare and high-risk traffic events. Our framework, Adv-BMT, benefits the development and testing of safer autonomous agents by exposing failure cases under challenging interactions. However, generating adversarial scenarios may potentially raise concerns about potential misuse, such as crafting unrealistic or malicious simulations. To address this, our approach is designed for research and evaluation within closed simulation environments. We encourage responsible usages of our model and encourage integrating them into safety validation pipelines with appropriate regulations.

### v2

@@
 However, such testing is limited by the scarcity of long-tailed, safety-critical scenarios in existing datasets collected in the real world.
-To tackle the data issue, we propose the Adv-BMT framework, which augments real-world scenarios with diverse and realistic adversarial interactions.
+To tackle the data issue, we propose the Adv-BMT framework, which augments real-world scenarios with diverse and realistic adversarial traffic interactions.
 The core component of Adv-BMT is a bidirectional motion transformer (BMT) model to perform inverse traffic motion predictions, which takes agent information in the last time step of the scenario as input, and reconstruct the traffic in the inverse of chronological order until the initial time step.
@@
 Different from previous work, we do not need any collision data for pretraining, and are able to generate realistic and diverse collision interactions.
-Our experimental results validate the quality of generated collision scenarios by Adv-BMT: training in our augmented dataset would reduce episode collision rates by 20% compared to previous work.
-The code will be made available at https://metadriverse.github.io/adv-bmt/.
+Our experimental results validate the quality of generated collision scenarios by Adv-BMT: training in our augmented dataset would reduce episode collision rates by 20%.
+Demo and code are available at https://metadriverse.github.io/adv-bmt/.
 1 Introduction
-In recent years, autonomous driving (AD) agents have achieved unprecedented performance in urban environments dosovitskiy2017carlaopenurbandriving ; li2022metadrivecomposingdiversedriving ; wu2024metaurbanembodiedaisimulation .
+In recent years, autonomous driving (AD) agents have achieved unprecedented performance in simulations dosovitskiy2017carlaopenurbandriving ; li2022metadrivecomposingdiversedriving ; wu2024metaurbanembodiedaisimulation .
 However, handling corner traffic situations, especially collision scenarios, remains a major challenge.
-One cause is that safety-critical scenarios are missing from real-world driving datasets due to high costs and risks of data collections.
+A major cause is that safety-critical scenarios are missing from real-world driving datasets due to high costs and risks of data collections.
 Without enough collision training data, it is hard for the autonomous driving (AD) planners and prediction models to learn safe driving in challenging and risky scenarios.
-This motivates the need for simulating collision traffic that emulates real-world accidents.
-According to rempe2022generatingusefulaccidentpronedriving , useful safety-critical scenarios should be realistic and diverse, as well as challenging.
-Previous works rempe2022generatingusefulaccidentpronedriving ; zhang2023catclosedloopadversarialtraining ; stoler2025sealsafeautonomousdriving ; ransiek2024goosegoalconditionedreinforcementlearning leveraged learned real-world traffic priors to constraint adversarial trajectories, and optimized predictions on collision-encouraging objectives.
-However, from our observations and evaluations, generated behaviors from these methods have a lack of diversity and a lower collision success rates (attacking efficiency).
-To improve upon previous work, we propose the Adv-BMT framework for realistic and diverse collision augmentations from real-world driving data.
-Adv-BMT samples various collision situations from the input driving log, and then reconstruct complete adversarial trajectories accordingly.
-This design enhances the diversity of predicted collision interactions, since now it augments different collision behaviors for each driving log.
-Besides, there is no need to include collision data in trainings.
-We further implement a rule-based rejection sampling mechanism to filter out candidate trajectories from unsatisfactory adversarial initializations, and maintain realistic collision interactions.
-In the core of Adv-BMT is the bidirectional motion transformer (BMT) model.
-Similar to existing multi-agent simulators and prediction models wu2024smartscalablemultiagentrealtime ; seff2023motionlmmultiagentmotionforecasting ; zhou2024behaviorgptsmartagentsimulation , BMT tokenizes continuous agent trajectories into discrete actions, and performs next-token predictions in an auto-regressive manner.
-Different from existing models, BMT learns to predict both future and history motion for all agents, conditioning on one-step current states.
-This is achieved by processing two sets of motion tokens from the input scenario, with one set of tokens used for forward prediction (future motion) and one for reverse prediction (history motion).
-As a summary, Adv-BMT is a two-staged pipeline: first, it initializes diverse collision frames between an new adversary agent (ADV) and ego vehicle; then, it reconstructs the adversarial trajectories via BMT’s reverse predictions.
-In the output, ADV is added to the original scenario, maintaining realistic interactions with surrounding traffic.
+This motivates the need for simulating different real-world accidents.
+To generate realistic collision trajectories, previous works rempe2022generatingusefulaccidentpronedriving ; zhang2023catclosedloopadversarialtraining ; stoler2025sealsafeautonomousdriving leveraged learned real-world traffic priors, and optimized predictions on collision-encouraging objectives.
+However, our evaluations of these baselines reveal that the generated behaviors are insufficiently diverse and yield a low collision generation rate.
+Figure 1: Overview of Adv-BMT.
+The framework mainly consists of three steps: (1) it adds an adversarial agent (ADV) with a sampled collision state into the current scenario; (2) it predicts reversely for the adversarial trajectory; (3) it performs rule-based checks and rejects physically implausible ones.
+We tackle this challenge by rethinking the motion prediction architecture itself.
+We introduce the Bidirectional Motion Transformer (BMT), a new model that learns to predict both future and history trajectories for all agents, conditioned on their current states.
+Similar to recent autoregressive traffic forecasting models wu2024smartscalablemultiagentrealtime ; seff2023motionlmmultiagentmotionforecasting ; zhou2024behaviorgptsmartagentsimulation , BMT tokenizes continuous trajectories into discrete control actions.
+Distinct from prior work, BMT employs a temporally reversible tokenization scheme that enables unified forward (future) and reverse (history) motion prediction within the same framework.
+We utilize BMT model into our Adv-BMT framework for realistic and diverse collision traffic generations from real-world driving data.
+While existing works follow a standard paradigm, which selects a convenient neighbor agent and modifies the behavior to attack the ego agent, Adv-BMT inserts new agents (ADV) with diverse collision interactions, maintaining realistic interaction with other traffic agents.
+In short, Adv-BMT is a three-staged pipeline: first, it initializes diverse collision frames between a new adversary agent (ADV) and ego vehicle; then, it reconstructs the adversarial trajectories via BMT’s reverse predictions; finally, it conducts rule-based checks and rejects trajectories with physically implausible collision initializations.
 An overview of our framework is illustrated in Fig.
 1.
+It is worth noting that that there is no collision data included in our model training.
+Adv-BMT is designed with multiple generation modes to support varying levels of agent interactions and realism.
+By default, traffic agents follow their recorded trajectories to preserve consistency with the real-world driving log, while the adversarial agent is generated to interact within this fixed context.
+This enables targeted scenario editing while maintaining overall scene plausibility.
+To address limitations in interactivity, Adv-BMT also supports a closed-loop reverse prediction mode, in which all agents are jointly predicted to generate a fully reactive traffic scenario, without teacher-forcing any agent behaviors during predictions.
+Additionally, an optional forward refinement step allows traffic agents to respond to the newly introduced adversary, enabling a more interactive and dynamically consistent outcome.
+Together, these modes allow Adv-BMT to balance realism, controllability, and diversities for different use cases.
 We summarize our contributions as follows: (1) We introduce the bidirectional motion transformer with temporally reversible motion tokenizations; (2) We develop Adv-BMT for realistic and diverse safety-critical traffic simulations; (3) We leverage Adv-BMT in a closed-loop setting to dynamically create challenging environments for reinforcement learning agents.
 2 Related Work
-Driving Motion Predictions and Simulations
-Recent work have utilized auto-regressive sequence modeling on discretized motion representations via transformer-based models.
-In previous works, Trajeglish philion2024trajeglishtrafficmodelingnexttoken discretizes motion representations using a k-disk based tokenization scheme to represent position and angle differences for relative movements; but the interaction performance is sensitive to agent ordering, and agents decoded first might not be as reactive to the rest of the predicted agents.
+Driving Motion Prediction.
+The task of motion prediction focuses on forecasting the future trajectories of traffic participants conditioned on their initial map context and agent states.
+Recent work uses transformer models to autoregressive sequence modeling.
+A set of work utilizes discretized motion tokenizations to perform next token predictions: Trajeglish philion2024trajeglishtrafficmodelingnexttoken discretizes motion representations using a k-disk-based tokenization scheme to represent position and angle differences for relative movements.
 MTR++ shi2024mtrmultiagentmotionprediction directly represents motion on continuous space in Gaussian mixture distributions.
-The design is efficient for multi-modal generations but prediction diversities across modes rely on the learned intention queries.
-MotionLM seff2023motionlmmultiagentmotionforecasting and SMART wu2024smartscalablemultiagentrealtime model motion actions that represents trajectory deltas on decoder-only transformers.
-BehaviorGPT zhou2024behaviorgptsmartagentsimulation performs next-patch predictions with future motion chunks, instead of single-step predictions, and achieves higher inference efficiency and reduces compounding errors.
-Safety-Critical Scenario Generations
-Previous works leverage learned traffic prior to generate realistic safety-critical interactions.
-STRIVE rempe2022generatingusefulaccidentpronedriving enhances diversity through modeling traffic a latent space using variational autoencoders but requires computationally expensive per-scenario optimizations.
-CAT zhang2023catclosedloopadversarialtraining utilizes the augmentation framework into MetaDrive li2022metadrivecomposingdiversedriving simulator as a closed-loop reinforcement learning environment.
-Upon CAT, SEAL stoler2025sealsafeautonomousdriving uses a skill-based adversarial policy for human-like driving guided by collision-related objectives.
-MixSim Suo_2023_CVPR generates realistic reactions by exploring future goals and re-simulates the inferred trajectories using a trained policy.
-GOOSE ransiek2024goosegoalconditionedreinforcementlearning uses a goal-conditioned reinforcement learning approach, which parameterizes trajectories and goal constraints to generate scenario-level, safety-critical interactions.
+MotionLM seff2023motionlmmultiagentmotionforecasting models trajectory deltas.
+SMART builds a k-disk–clustered vocabulary of motion tokens, containing coordinates, heading, and shape.
+BehaviorGPT zhou2024behaviorgptsmartagentsimulation performs next-patch predictions with future motion chunks, instead of single-step predictions.
+The BMT model constructs two sets of motion tokens based on inter-frame accelerations, enabling both forward prediction of future motions and inverse reconstruction of past motions.
+Another set of works Jiang_2023_CVPR ; zhong2022guidedconditionaldiffusioncontrollable ; zhong2023languageguidedtrafficsimulationscenelevel ; 10610324 ; 11127600 ; NEURIPS2023_d95cb79a ; Xu_Petiushko_Zhao_Li_2025 ; 10801623 leverage diffusion models for motion predictions.
+Safety-Critical Traffic Scenario Generation.
+To generate collision traffic, a classic line of previous works use a two-staged method: first uses a traffic prior to generate realistic agent trajectories, then use collision-sensitive objectives for trajectory optimization or candidate selections.
+STRIVE rempe2022generatingusefulaccidentpronedriving models traffic by a variational autoencoders but requires computationally expensive per-scenario optimizations.
+CAT zhang2023catclosedloopadversarialtraining leverages transformer motion decoder gu2021densetntendtoendtrajectoryprediction and selects the collision trajectories and simulates in MetaDrive li2022metadrivecomposingdiversedriving for closed-loop adversarial environments.
+SEAL stoler2025sealsafeautonomousdriving uses a skill-based adversarial policy with collision-related objectives.
+SafeSim chang2024safesimsafetycriticalclosedlooptraffic proposes a diffusion model with a test-time collision-sensitive guidance loss to control the collision type and adversarial agent selections.
+Another set of work such as ransiek2024goosegoalconditionedreinforcementlearning use a reinforcement learning (RL) based approach, which parametrizes trajectories and goal constraints to generate safety-critical interactions.
 AdvSim wang2023advsimgeneratingsafetycriticalscenarios directly perturbs actor trajectories using a kinematic model and optimizes via a black-box adversarial loss.
-Recent work ding2020multimodalsafetycriticalscenariosgeneration uses a conditional normalizing flow to model the distribution of real-world safety-critical trajectories.
-SafeSim chang2024safesimsafetycriticalclosedlooptraffic proposes a diffusion model with a test-time collision-sensitive guidance loss to control the collision type and adversarial agent selections; however its generation quality relies on initial collections of collision data.
+Another line of recent work such as ding2020multimodalsafetycriticalscenariosgeneration uses a conditional normalizing flow to model the distribution of real-world safety-critical trajectories.
 CrashAgent li2025crashagentcrashscenariogeneration and LCTGen tan2023languageconditionedtrafficgeneration leverage free-form texts as inputs and extract embeddings to parameterize scene initializations and agent driving directions.
-Figure 1: Overview of Adv-BMT framework.
-From the input scenario, Adv-BMT first samples diverse collision states between ego agent and a new adversarial agent, then reconstruct complete adversarial trajectories via BMT’s reverse motion predictions.
-A rule-based reject sampling mechanism is used for filtering unsatisfactory collision initializations.
-3 Methodology
-In this section, we formulate forward and reverse motion predictions. We then introduce bidirectional motion transformer (BMT) model. Last but not the least, we demonstrate details of Adv-BMT.
-3.1 Problem Formulation
-Consider a traffic scenario with maximumly $N$ agents and time horizon $T$.
-The trajectory of agent $i$ is denoted by $\boldsymbol{\tau}^{i}=\{{\tau}^{i}_{0},{\tau}^{i}_{1},\dots,{\tau}^{i}_{T}\}$, where each state $\tau^{i}_{t}\in\mathbb{R}^{d}$ includes position, velocity, and heading at time $t$.
-We define a prediction direction indicator $D\in\{\text{Forward},\text{Reverse}\}$, for future or past motion prediction within $T$ time steps.
-For each scenario we construct a sequence of motion tokens:
-$\mathbf{Z}^{i}=\{z^{i}_{1},\dots,z^{i}_{T}\},\quad z^{i}_{k}=\begin{cases}\phi(\tau^{i}_{k},\tau^{i}_{k-1}),&D=\text{Forward}\\ \phi(\tau^{i}_{T-k},\tau^{i}_{T-k+1}),&D=\text{Reverse}\end{cases}$ (1)
-where $\phi(\cdot)$ denotes the motion tokenization function which we discuss in section 3.2.
-The goal is to model the distribution of agent motion tokens at the $k$-th step, conditioned on the predicted motion token sequence, current state, and scene context $\mathcal{M}$:
-$P(z^{1}_{k},z^{2}_{k},\dots,z^{N}_{k}\mid\mathbf{z}^{1}_{1:k-1},\mathbf{z}^{2}_{1:k-1},\dots,\mathbf{z}^{N}_{1:k-1},\mathcal{M},D)$ (2)
-where $\mathbf{Z}^{i}_{0:k-1}$ denotes history token sequence for agent $i$.
-The decoder autoregressively selects the next action token $z^{i}_{k+1}$ by attending to past tokens and scene token embedding $\mathcal{M}$.
-Each selected action in forward prediction maps to a position in future time-step, similar to a standard motion predictor; selected token sequence in reverse prediction can be reconstructed as history trajectories.
-Together, BMT is able to perform both future and past motion predictions as shown in Fig.
-4.
-Figure 2: BMT model architecture.
-3.2 Bi-direction Motion Tokenizations
-Token space
+Different from previous approaches, Adv-BMT generates collision scenarios in three steps: first samples a collision state, then conduct reverse predictions, finally forcast traffic agents accordingly.
+3 Method
+Classical motion prediction models forecast future trajectories based on the current states of traffic agents.
+Building on this foundation, we propose the reverse motion prediction problem.
+To address both tasks within a unified framework, we introduce the Bidirectional Motion Transformer (BMT) model, which is able to perform both tasks.
+Finally, we present the Adv-BMT framework and describe how it leverages BMT model as the core to generate realistic and diverse safety-critical interactions.
+Figure 2: Bidirectional predictions on the ego agent (red). BMT supports predictions for future motions (left) and historical motions (right) jointly for all prediction agents.
+3.1 Bidirectional Motion Prediction
+We first introduce the bidirectional motion prediction task, shown in Fig.
+2.
+Consider a traffic scenario with at most $N$ agents and a prediction horizon of $T$ steps.
+The trajectory of agent $i$ is represented as $\boldsymbol{\tau}^{i}=\{\tau^{i}_{0},\tau^{i}_{1},\dots,\tau^{i}_{T}\}$, where each state $\tau^{i}_{t}\in\mathbb{R}^{d}$ encodes its position, velocity, and heading at time $t$.
+We introduce a prediction direction indicator $D\in\{\text{Forward},\text{Reverse}\}$, to specify whether the model predicts future or past motion over the horizon.
+For each predicted traffic agent, we construct a sequence of motion tokens $\mathbf{Z}^{i}=\{z^{i}_{1},\dots,z^{i}_{T}\}$, by applying the motion tokenization function $\phi(\cdot)$ between consecutive states.
+In the forward setting, the tokens are derived in chronological order, from $\tau^{i}_{0}$ to $\tau^{i}_{T}$.
+In the reverse setting, the temporal order of the trajectory is inverted, and the tokens are generated by applying $\phi(\cdot)$ backward from the current state toward the initial state.
+This formulation yields a bidirectional token sequence that enables BMT to model both forward and reverse motion dynamics in a unified token space.
+The tokenization function $\phi(\cdot)$ is further detailed in Section 3.2.
+3.2 Bidirectional Motion Transformer (BMT)
+Token Space.
 BMT’s bidirectional motion tokens are derived from a simplified bicycle dynamics model.
@@
 We uniformly quantize the control space of accelerations $a\in[-a_{\max},a_{\max}]$ and yaw rates $\delta\in[-\delta_{\max},\delta_{\max}]$ into $K$ bins each, yielding a total of $K^{2}$ discrete motion tokens, where $a_{\max}=10m/s$, $\delta_{\max}=\frac{\pi}{2}$, and $K=33$.
-Control token simulation
-Given a motion control token, BMT simulates it and reconstructs to a new agent state using midpoint integration:
-$\text{Recons}(z,\tau,\Delta t)=\left(x+\bar{v}\cdot\cos(\bar{\theta})\cdot\Delta t,\;y+\bar{v}\cdot\sin(\bar{\theta})\cdot\Delta t,\;\theta+\omega\cdot\Delta t,\;v+a\cdot\Delta t\right)$ (3)
-where $\tau=(x,y,\theta,v)$ denotes the current agent state with position $(x,y)$, heading $\theta$, and speed $v$; $z=(a,\omega)$ is the motion token containing acceleration $a$ and yaw rate $\omega$; $\Delta t$ is the time interval; $\bar{v}=v+\frac{1}{2}\cdot a\cdot\Delta t$ is the midpoint speed; and $\bar{\theta}=\theta+\frac{1}{2}\cdot\omega\cdot\Delta t$ is the midpoint heading.
-Ground-truth tokens
-For obtaining ground truth motion tokens during trainings, we simulate all $K^{2}$ candidate motion tokens starting from each agent state at time $t$, then select the token whose predicted next state $\hat{\tau}_{t+1}$ best matches the ground truth $\tau_{t+1}$.
-This process maps the continuous trajectory into a sequence of discrete motion tokens $\{z_{0},z_{1},\dots,z_{T}\}$ for each predicted agent.
-Forward and Reverse predictions share the same motion token space but differ in token interpretations.
-In reverse predictions, we invert the temporal order of the trajectory and treat the final state (e.g., collision initializations from Adv-BMT) as the initial; at each time step, we simulate each motion token (by inverting the time step, $\Delta t\to-\Delta t$), and find the token that best reconstructs the previous state $\tau_{t}$ from $\tau_{t+1}$.
-This generates a motion token sequence in reverse time direction $\{\tilde{z}_{0},\tilde{z}_{1},\dots,\tilde{z}_{T}\}$, where $\tilde{z}_{0}$ corresponds to the final state in forward time.
-3.3 Token Matching with minimized Contour Error
-For either reverse or forward predictions, we build motion tokens step-by-step in the order of prediction time, starting from the first prediction step and reconstructing till the last prediction step.
-With a 0.5-second time interval, a search is conducted from all candidate token vocabularies, and selects the best candidate that minimizes the distance between the ground-truth and reconstructed contours for the $i$-th candidate action for agent $n$:
-$a_{n}^{*}=\arg\min_{i}\left(\mathcal{E}_{n}^{(i)}=\frac{1}{P}\sum_{p=1}^{P}\left\|\mathbf{c}_{n}^{(i,p)}-\hat{\mathbf{c}}_{n}^{(p)}\right\|_{2}\right)\quad\text{where }0\leq i<K^{2}$ (4)
-where $\mathbf{c}_{n}^{(i,p)}\in\mathbb{R}^{2}$ and $\hat{\mathbf{c}}_{n}^{(p)}\in\mathbb{R}^{2}$ represent the $p$-th polygon corner of the reconstructed and ground-truth contours, each with totally $P$ contour points; $a_{n}^{*}$ is the index of the selected candidate token.
-3.4 Bidirectional Motion Transformer (BMT)
-The BMT architecture and training process is shown in Fig. 2, along with ground-truth token matching for forward predictions or reverse predictions.
-Architecture
+Motion Token Reconstruction.
+A key step in BMT is to reconstruct continuous trajectories from discrete motion tokens, bridging between physical motion and the token vocabulary.
+Each token encodes a low-level control command defined by an acceleration $a$ and a yaw rate $\omega$.
+Starting from the current agent state $\tau=(x,y,\theta,v)$, the token is applied over a time interval $\Delta t$ to update the agent’s position, heading, and velocity.
+BMT adapts midpoint integration to propagate the state, which provides a stable and realistic approximation of vehicle dynamics.
+Intuitively, the speed and heading are updated using their midpoint values, ensuring that the reconstructed trajectory remains smooth and physically consistent.
+This reconstruction procedure allows BMT to decode token sequences into continuous motion trajectories in both forward and reverse prediction settings.
+Model Architecture.
+The BMT architecture overview is shown in Fig.
+3.
 BMT has a scene encoding component used to obtain embeddings for scenario contexts with separate embeddings for map polylines, traffic lights, and agent initial states.
 Then, we use Fourier-encoded edge features tancik2020fourierfeaturesletnetworks to represent the spatial and directional information between these encoded entities, which are then passed to the transformer encoder with self-attention layers for the relational embeddings.
+Figure 3: BMT architecture.
+BMT consists of a scene encoder and a GPT-style motion decoder.
+It employs two sets of motion tokens for forward and reverse predictions to generate the next-step token for each agent.
+All predictions are conditioned only on the map information and the one-step current state of all predicted agents.
 The prediction decoder predicts subsequent motion tokens in an autoregressive manner, with only initial agent information for the first frame, along with the scene embedding obtained from the Scene Encoder.
@@
 The output agent motion embeddings are mapped to the vocabulary of discretized motion tokens through MLPs.
-Training
-In training, BMT predicts the sequence of motion tokens and use the cross-entropy objective to match the joint action distribution of the observed behaviors in the dataset during training:
-$\max_{\pi_{\theta}}\;\mathbb{E}_{\mathcal{D}}\left[\sum_{t=1}^{T}\sum_{i}^{N}\log\pi_{\theta}\left(z^{i}_{t}\mid\mathbf{z}^{i}_{1:k-1},\mathcal{M},D\right)\right]$ (5)
-During inference, BMT generate motion tokens autoregressively, and uses nucleus sampling (top-$p$ sampling) to enhance diversity and maintain plausibility.
-To prevent the exposure bias during auto-regressive inference, BMT rolls out upon chosen motion tokens instead of the ground-truth token sequence.
-3.5 Adv-BMT
-The overview of Adv-BMT is illustrated in Fig.1.
-The framework mainly consists of three steps: (1) it adds an adversarial agent (ADV) with a specified collision state into the current scenario; (2) it predicts reversely for the adversarial trajectory; (3) it performs rule-based checks and reject physically implausible ones.
-Diverse Adversarial Initializations
-While existing works select a convenient neighbor agent and modify the behavior to attack the ego agent, Adv-BMT inserts new agents (ADV) with diverse collision initializations for different opponent interactions.
-We visualize an example result in 4.
+More details can be found in the Appendix.
+Training.
+BMT is trained to learn a policy that reproduces the distribution of real-world driving behaviors.
+At each step, the model predicts a discrete motion token for every agent based on its past tokens, current state, and the scene context.
+To align these predictions with the ground-truth behaviors in the dataset, we minimize a cross-entropy loss between the predicted and observed token distributions:
+$\mathcal{L}_{\text{train}}=-\;\mathbb{E}_{\mathcal{D}}\left[\sum_{t=1}^{T}\sum_{i=1}^{N}\log\pi_{\theta}\left(z^{i}_{t}\mid\mathbf{z}^{i}_{1:t-1},\mathcal{M}\right)\right],$ (1)
+where $\pi_{\theta}$ is the token prediction policy parameterized by $\theta$.
+Intuitively, this objective encourages the model to assign high probability to motion tokens that correspond to real trajectories, thereby capturing the joint distribution of multi-agent actions in traffic.
+During inference, BMT generates motion tokens autoregressively, sampling one token at a time conditioned on its previous predictions.
+We apply nucleus (top-$p$) sampling to promote behavioral diversity while remaining faithful to the learned distribution.
+To mitigate exposure bias, the model rolls out on its own sampled tokens rather than the ground-truth sequence, ensuring consistency between training and inference.
+3.3 Adv-BMT for Safety-critical Interaction
+The BMT model serves as the core that captures the distribution of realistic multi-agent motions in Adv-BMT framework.
+Building upon this foundation, we develop Adv-BMT, an adversarial scenario generation framework designed to generate realistic and safety-critical traffic interactions.
+The overview of Adv-BMT framework is illustrated in Fig.1.
+From the input scenario, Adv-BMT first samples diverse collision states between ego agent and a new adversarial agent (ADV), then reconstructs complete trajectories via reverse motion predictions.
+A rule-based reject sampling mechanism is used for selecting collision initializations.
+Figure 4: Diverse collision headings generated by Adv-BMT from the same driving log input, visualized from a third-person perspective.
+Figure 5: Diverse collision timings generated by Adv-BMT from the same driving log input, visualized in a bird’s-eye view.
+Diverse Adversarial Initialization.
+While existing works select a convenient neighbor agent and utilize its initial information, Adv-BMT inserts new agents (ADV) with diverse collision initializations for different opponent interactions.
+We visualize an example result in 5.
 Formally, we define the collision state as to include the position, time, velocity, and heading at the collision step.
-The collision time is sampled from the first second to the last timestep of the ego trajectory length.
+The collision time can be varied, sampled from the first second to the last time step of the ego trajectory length.
 Similarly, ADV’s collision headings are randomly sampled.
-ADV’s collision position can be calculated from collision heading and the ego vehicle’s collision position, with the ADV contour intersecting with the ego vehicle’s contour.
-Last but not least, collision velocity is also calculated from sampling a variable offset from ego vehicle’s speed at collision step.
-Reverse Predictions
-We use the sampled collision states as inputs of BMT for reverse predictions, and obtain ADV’s trajectory from predictions to augment the original scene with a new adv agent.
-During BMT inference, we teacher-force the ground-truth token sequence of ego agent and traffic agents, to keep the originality of the scenario as much as possible (since BMT could predict agents to driving on a different lane choice, which makes it a totally different scenario).
-By teacher-forcing traffic vehicle agents, BMT reconstructs the adversarial agent trajectory to fit in the input scenario.
-Reconstructing entirely new scenarios is out of the scope of our current work, but BMT is able to realize both normal and safety-critical generations through forward and reverse predictions.
-Rule-based Rejection Samplings
+ADV’s collision position can be calculated from collision heading and the ego vehicle’s collision position.
+Last but not least, collision velocity is calculated from sampling a offset from ego vehicle’s speed at collision step.
+Multi-agent Adversarial Interaction.
+A key advantage of Adv-BMT lies in its ability to generate realistic and flexible multi-agent interactions surrounding the adversarial agent (ADV).
+Starting from sampled collision initializations, the ADV is added into the scene together with existing traffic participants.
+BMT then performs reverse-time prediction to reconstruct plausible interaction histories that lead to the designated collision state.
+To balance realism, controllability, and scene reactivity, Adv-BMT provides several generation modes with different levels of multi-agent coupling:
+  1. 1.
+    Adv-BMT with replayed traffic agents: In this mode, traffic agents are teacher-forced to follow their original recorded trajectories, while only the adversarial agent’s history is predicted in reverse time.
+This preserves the original traffic flow and ensures high scene fidelity.
+Teacher-forcing prevents unintended deviations (e.g., traffic agents switching lanes or diverging from the original trajectory) that may occur if the model predicts all agents in the closed-loop manner.
+As a result, BMT reconstructs the adversarial agent’s trajectory to fit into the original scenario.
+  2. 2.
+    Adv-BMT with closed-loop reverse prediction: This mode removes teacher-forcing and jointly predicts the trajectories of all agents, including both traffic participants and the adversarial agent, in a single reverse prediction pass.
+By allowing all agents to evolve backward in time simultaneously, the scene develops coherently as a fully interactive scenario.
+This setting enables rich multi-agent dynamics and can produce fully synthesized safety-critical scenarios.
+  3. 3.
+    Adv-BMT with forward refinement: This hybrid mode first applies teacher-forced reverse prediction to generate the adversarial agent, then runs an additional forward prediction pass to enable traffic agents to react to its presence.
+This introduces interaction and reactivity without discarding the structure of the original recorded scene, offering a middle ground between strict replay and fully closed-loop generation.
+These complementary modes provide a controllable trade-off between scene realism, interactivity, and controllability, enabling users to select the most suitable configuration for different evaluation objectives.
+Rule-based Rejection Sampling.
 We design Adv-BMT to have diverse initializations, which do not guarantee the realism of the collision outcomes.
-To address this issue, we implement a rule-based rejection mechanism for ADV candidates.
-We first measure the driving distance and average speed of ADV candidates; if it moves too short and mostly wanders at the designated position waiting for the ego vehicle, then it is considered invalid.
+To address this issue, we implement a rule-based filtering rejection mechanism for ADV candidates.
+Specifically, we first measure the driving distance and average speed of ADV candidates; if it moves too short and mostly wanders at the designated position waiting for the ego vehicle, then it is considered invalid.
 Meanwhile, we check the max curvature (the rate of change for heading): given a candidate ADV trajectory, we compute the curvature constraint using: $\kappa_{t}={\Delta\theta_{t}}/{\Delta s_{t}},$ where $\Delta\theta_{t}$ is the absolute heading change between time steps, and $\Delta s_{t}$ is the displacement between consecutive positions.
 Adv-BMT rejects a prediction if $\max_{t}(\kappa_{t})>\kappa_{\text{threshold}}$, where $\kappa_{\text{threshold}}=0.8$ is a predefined curvature limit that we found useful.
-We check the curvature to ensure it does not exceed our defined threshold, and reject candidate ADVs with unrealistic changes in headings.
-With our straightforward rule-based rejection sampling mechanism, we are able to maintain only realistic collision events between the ADV and the ego vehicle.
-Figure 3: BMT predicts both future and history agent trajectories.
-Figure 4: Adv-BMT generates diverse collision scenarios.
-Forward refinement for Reactive Traffic
-In real-world collision events, neighbor vehicles would be affected by the collision vehicles and would have reactions like sudden lane changes and brakings.
-With this concern, we implement Bi-directional Adv-BMT, with a forward refinement to re-simulate traffic agents upon predicted ADV trajectory.
-In forward refinement, all traffic agents would be predicted and modified in the output scenario, with teacher-forcing motions of collision participants (ego and ADV agents).
-In experiments, we refer to it as Adv-BMT (Bi) to indicate bi-directions.
+We enforce a curvature constraint to ensure that the trajectory remains within the predefined threshold, rejecting ADV candidates that exhibit unrealistically sharp turning behavior.
+With our straightforward rule-based rejection sampling mechanism, we are able to maintain realistic collision events between the ADV and the ego vehicle.
 4 Experiments
-In section 4.1, we evaluate BMT model on motion prediction performance.
-In section 4.2, we compare Adv-BMT with baseline methods.
-In section 4.3 we train a reinforcement learning agent in different safety-critical augmented datasets.
-For augmentations, we use ground-truth traffic data from the Waymo Open Motion Datasets (WOMD) ettinger2021largescaleinteractivemotion .
-4.1 Realism of Adv-BMT
-Settings
-We measure BMT on four different settings: Forward prediction with or without agent history, Reverse prediction without agent history, and bi-directional predictions.
-We measure on open-loop scenario generation metrics with 6 prediction modes on each scenario: scenario final displacement errors (SFDE) which measures the average displacement error at the final step on the best prediction mode, and scenario average displacement errors (SADE) which measures the averages according to all time steps on the best mode.
-We also measure Final Displacement Diversity (FDD) and average displacement diversity (ADD) for prediction diversities across different prediction modes: FDD measures the spread of final positions and ADD measures trajectory variations over time.
-Finally, we measure the overall collision rates for realistic interactions.
-The results are presented in Table 1.
-Table 1: Evaluation for BMT on scenario realism on different prediction modes.
-Context SFDE $\downarrow$ SADE $\downarrow$ minSFDE $\downarrow$ minSADE $\downarrow$ FDD $\uparrow$ ADD $\uparrow$ Collision
-Forward 1s 3.043 1.019 1.908 0.693 9.241 3.240 0.026
-Forward 0s 3.180 1.297 2.144 0.927 10.263 4.177 0.028
-Reverse 0s 3.411 1.380 2.158 0.975 8.234 2.913 0.031
-Bidirections 0s 3.059 2.111 1.831 1.536 7.618 6.270 0.035
-Analysis
-Results from Table 1 indicate that BMT performs realistic scenario generations in both forward predictions and reverse predictions.
-Furthermore, by conditioning on a 1-second history, the forward prediction achieves better SADE and SFDE scores, indicating better temporal consistencies.
-Due to design of Adv-BMT with agent initializations at the last step, it is impossible for reverse predictions to access agent histories.
-Compared to single-pass prediction, bi-directional prediction exhibit lower prediction accuracy in ADE metrics, indicating a harm in the realism of generations.
-In terms of prediction diversities, Forward prediction exhibits higher FDD and ADD scores than reverse prediction.
-This indicates that predicting histories is more constrained, but predicting future motions encourages more explorations for variant directions.
-This further validates the necessity of Adv-BMT’s diverse collision initialization design to force different collision outcomes.
-Despite these differences, the overall collision rate remains reasonable and comparable for all prediction modes, and indicates that both prediction modes effectively generate realistic interactions.
-Figure 5: Comparison among adversarial behaviors of different methods.
-4.2 Safety-critical Evaluations
-In following experiments, We compare Adv-BMT with three baseline methods for safety-critical augmentations, namely STRIVE rempe2022generatingusefulaccidentpronedriving , CAT zhang2023catclosedloopadversarialtraining , and SEAL stoler2025sealsafeautonomousdriving .
-We evaluate the quality and efficiency of generated collisions among all methods.
-We measure the adversarial attack success rate (SDC-ADV collision rate), collision rates between adversary agents and traffic agents (ADV-BV Collision rate), indicating interaction realism.
-Additionally, we evaluate the difference between the predicted distribution and the ground-truth distribution on velocities (VEL JSD) and accelerations (ACC JSD).
-These two metrics also indicate the realism of generated adversarial agent motions.
-Settings
-We randomly select 500 WOMD scenarios, each as a base environment for adversarial augmentations.
-For each scenario, each method predicts six different adversarial trajectories.
-To be consistent with baseline methods in this evaluation only, we force Adv-BMT to choose an existing neighbor agent and modify its behavior as the predicted adversary.
-Table 2: Evaluation for Safety-critical Scenarios
-[…]
-CAT zhang2023catclosedloopadversarialtraining 0.470 0.271 0.137 0.480 0.0 0.0
-SEAL stoler2025sealsafeautonomousdriving 0.582 0.261 0.239 0.561 0.0 0.0
-STRIVE rempe2022generatingusefulaccidentpronedriving 0.534 0.358 0.117 0.432 0.004 0.001
-Adv-BMT 1.000 0.391 0.142 0.435 0 2.344
-Adv-BMT (Bi) 0.997 0.373 0.188 0.534 0.038 2.257
-Quantitative Results
-Table 2 shows the quantitative results for scenario generation metrics of all methods.
-Two settings of Adv-BMT outperform baselines in terms of adversarial attack success rate.
+We first assess the BMT model on its ability to generate realistic and diverse traffic behaviors in Section 4.1.
+We then evaluate quality of Adv-BMT scenarios compared to three baseline methods in Section 4.2.
+Furthermore, in Section 4.3 we evaluate the utility for the downstream learning task, and assess whether training a reinforcement learning (RL) planner in Adv-BMT scenarios lead to improved performance and robustness compared to log-replay traffic flows.
+Dataset.
+All experiments use driving data from the Waymo Open Motion Dataset (WOMD) ettinger2021largescaleinteractivemotion with formats managed by ScenarioNet li2023scenarionetopensourceplatformlargescale .
+WOMD contains 10Hz scenarios, each with 1 second of history and 8 seconds of future trajectories.
+Each scenario includes up to 128 traffic agents including vehicles, cyclists, pedestrians along with high-definition maps.
+To reduce computational cost, we downsample each scenario to 2Hz, yielding 19 prediction steps.
+We randomly select 500 scenes for both open-loop evaluation and RL training.
+We use 6 prediction modes for each scene during open-loop evaluations in Section 4.1 and Section 4.2.
+Metrics.
+To evaluate the realism of BMT predictions, we adopt standard open-loop prediction accuracy metrics: Scenario Final Displacement Error (SFDE) and Scenario Average Displacement Error (SADE).
+For each, we report results on both the best prediction mode and the average over all six prediction modes.
+We also measure the average number of agent collisions to capture interaction realism.
+To assess the diversity of generated trajectories, we report Final Displacement Diversity (FDD), Starting Displacement Diversity (SDD), and Average Displacement Diversity (ADD), which quantify the spread of predicted positions across all prediction modes.
+We also compute the Jensen–Shannon Divergence (JSD) values over velocity, acceleration, and time-to-collision (TTC) distributions between generated and ground-truth trajectories.
+JSD measures the similarity between two probability distributions, with lower values indicating closer alignment between generated and real-world behaviors.
+4.1 Evaluation of BMT Predictions
+Table 1: We evaluate BMT under three types of prediction tasks in a closed-loop setting for all traffic agents, without replaying any agents.
+In the bidirectional mode, forward prediction is performed based on the reverse prediction results.
+The number of prediction modes is set to six.
+(a) Realism of BMT predictions.
+Method $_{\text{avg}} SFDE$$_{\text{min}}$ $_{\text{avg}} SADE$$_{\text{min}}$ $_{\text{min}} VehColl$$_{\text{avg}}$ $_{\text{velocity}} JSD$$_{\text{TTC}}$
+Reverse 3.92 2.70 5.53 5.11 0.03 0.05 0.17 0.02
+Forward 3.52 2.35 2.39 1.98 0.03 0.05 0.23 0.30
+Reverse + Forward 3.29 2.01 2.38 1.74 0.03 0.06 0.22 0.35
+(b) Diversity of BMT predictions.
+Method SDD FDD ADD
+Reverse 8.35 - 3.13
+Forward - 10.78 4.40
+Reverse + Forward 8.49 12.78 7.00
+Results from Table 1(a) indicates that BMT performs realistic scenario generations in forward, backward, and bidirectional predictions.
+Reverse predictions perform slightly worse than forward predictions in realism metrics.
+Compared to single-pass prediction, bidirectional prediction exhibit higher prediction accuracy.
+Despite these differences, the overall collision rate remains reasonable and comparable for all prediction tasks, which indicates that BMT effectively generate realistic traffic interactions.
+The results in Table 1(b) show that bidirectional prediction enhances the diversity of generated agent behaviors across all displacement diversity metrics.
+Forward prediction exhibits higher FDD and ADD scores than reverse prediction.
+This indicates that predicting future motions encourages more explorations for variant directions.
+4.2 Evaluation of Adv-BMT Scenarios
+In Table 2, we additionally use the adversarial attack success rate (i.e., the collision rate between adversary and the ego), ADV-Traffic Collision Rate (i.e.,collision rates between adversary agents and traffic agents), and the average Agent Collision (i.e., average traffic agent collision rate) to indicate interaction realism.
+Two settings of Adv-BMT outperform baselines in both realism and diversity.
 This validates Adv-BMT’s design of collision initialization + reverse predictions, which couldn’t be achieved by other methods.
 While Adv-BMT generates highly adversarial behaviors, it also preserves diversity of traffic interactions compared to baselines.
-Note tha FDD and SDD metrics suggest that our baselines generate nearly the same adversarial trajectory on the given scene, while Adv-BMT is able to generate diverse behaviors at different prediction mode.
-The JSD metrics suggests slightly better realism on velocities and accelerations for baselines compared to Adv-BMT, showing the trade-off between diversity and velocity/acceleration realism.
-BMT model is able to achieve realistic motion predictions indicted by Waymo Open Sim Agent Challenge metrics montali2023waymoopensimagents , which we append in our appendix.
-Overall, our safety-critical evaluations indicate that Adv-BMT is an efficient framework for realistic, diverse, and safety-critical generations.
-Visualizations
-We simulate Adv-BMT scenarios in MetaDrive li2022metadrivecomposingdiversedriving simulators using both BEV and 3D rendering.
+FDD and SDD metrics suggest that baselines generate nearly the same adversarial trajectories on the given scene.
+The JSD metrics suggest Adv-BMT outperforms in realism metrics compared to baselines.
+BMT model is able to achieve realistic motion predictions indicted by Waymo Open Sim Agent Challenge metrics montali2023waymoopensimagents , which we add in Table 6 in our appendix.
+Furthermore, the results for Adv-BMT with filtering indicate that our rule-based filter does not harm diversity and at the same time enhances the realism metrics.
+Table 3 demonstrates the generation speed comparisons among all methods.
+Our evaluations validate Adv-BMT as an efficient framework for realistic, diverse, and safety-critical generation.
+Table 2: Diversity and realism of generated adversarial behaviors.
+We compare Adv-BMT scenarios with three baseline methods for safety-critical interactions, namely STRIVE rempe2022generatingusefulaccidentpronedriving , CAT zhang2023catclosedloopadversarialtraining , and SEAL stoler2025sealsafeautonomousdriving .
+To be consistent with baseline methods in this evaluation only, we force Adv-BMT to choose nearest neighboring agent and modify its behavior as the predicted adversary.
+[…]
+CAT 0.00 0.00 0.13 0.16 0.12 0.48 0.10 0.3
+SEAL 0.00 0.00 0.18 0.25 0.17 0.55 0.12 0.32
+STRIVE 0.01 0.00 0.11 0.22 0.12 0.38 0.09 0.32
+BMT_TF 1.22 0.49 0.10 0.17 0.13 1.00 0.13 0.22
+BMT_All 1.51 0.63 0.10 0.17 0.08 1.00 0.13 0.19
+BMT_TF + Filter 2.32 1.00 0.08 0.12 0.09 1.00 0.13 0.12
+BMT_All + Filter 1.97 0.83 0.13 0.14 0.08 1.00 0.12 0.09
+Visualization.
+We simulate Adv-BMT scenarios in MetaDrive li2022metadrivecomposingdiversedriving , rendering in both bird’s-eye view and third-person perspective.
 In Fig.
-4, we show diverse Adv-BMT adversarial behaviors in the collision directions on the same ego agent in an intersection scenario from WOMD.
-Adversarial agents (in red) come from different lanes and collide with the ego agent (in green) at different angles and facets.
-Besides, the generated adversarial agent follows the traffic rules and maintains realistic driving patterns.
-Apart from diverse collision directions, Adv-BMT generates collisions at diverse timings.
-Adv-BMT adversarial agents cover collisions with the ego before entering the intersection, during the middle of the intersection, and after turning to the main lane.
-With a limited number of real-world maps and logs, Adv-BMT is able to generate diverse collision behaviors and timings, making it suitable for AD testing and adversarial training.
-A visual comparison among Adv-BMT and baseline methods is shown in Fig.
-5, where other methods fail to generate an adversarial trajectory but Adv-BMT is able to insert a realistic ADV and triggers collisions .
-Generation Speeds
-Adv-BMT and CAT generate safety-critical scenarios significantly faster than other baselines.
-Adv-BMT generates one scenario in 1.02s compared to SEAL’s 2.36s and STRIVE’s 9.53s, achieving great efficiency gains.
-With details shown in Table 7 in Appendix.
+5, Fig.
+4, and Fig.
+9, we show diverse Adv-BMT adversarial behaviors in the collision directions on the same ego agent in several real-world traffic flows from WOMD.
+Generated adversarial agents follow traffic rules and maintain realistic driving patterns.
+Adv-BMT supports diverse adversarial agent types including vehicles, pedestrians, and cyclists, as shown in Fig.
+6.
+Adv-BMT makes full use of each driving log, which makes Adv-BMT suitable for AD testing and adversarial training.
+A visual comparison between Adv-BMT and baseline methods is shown in Fig.
+7.
+When a baseline fails to select an existing traffic agent that is convenient for a safety-critical attack, Adv-BMT is able to imagine a new agent at an appropriate position to perform adversarial attacks on the ego vehicle.
+Table 3: Generation speeds across methods. Among the four methods, CAT achieves the fastest generation speed, followed by Adv-BMT with a slightly lower speed. SEAL and STRIVE are comparatively slower.
+Method CAT SEAL STRIVE Adv-BMT
+500-avg (seconds) 0.80 2.36 9.53 1.02
 4.3 Adversarial Learning
-To validate the value of Adv-BMT scenarios in downstream autonomous driving (AD) tasks, we train a reinforcement learning agent within augmented scenarios with collision interactions generated by Adv-BMT and baseline methods.
-To determine the quality of augmented training scenarios, we measure both driving performance and safety performance of the learned AD agent compared to learning in original training set.
-In our experiment, we conduct two sets of trainings.
-(1) open-loop training: agent is trained in the fixed training set with adversarial scenarios generated based on ground-truth ego trajectories.
-(2) closed-loop training: an adaptive adversarial agent attacks on current ego agent based on its recent rollout trajectory records.
-The adaptive adversarial agent’s motion is generated by Adv-BMT and baseline.
-Results of both trainings are shown in table 3 and table 4 respectively.
-Settings
-The training set contains 500 real-world scenarios randomly selected from WOMD training set.
-Each waymo scenario is augmented with one collision scenario, so that the ratio between waymo and safety-critical scenes is 1:1.
-For each method, we augments the original dataset into a new training set.
-For the adaptive adversarial learning experiment, we implement an adaptive generator for Adv-BMT and CAT zhang2023catclosedloopadversarialtraining (discard other methods due to lower inference time).
-We train a Twin Delayed DDPG (TD3) agent for 1 million steps using 8 random seeds to ensure robustness.
-We conduct the training in the MetaDrive simulator and use the TD3 algorithm for reinforcement learning (hyper-parameters listed in Appendix).
-We measure the average reward, average cost, average route completion rate (Compl.), and average episode cost (cost sum) for driving performance measure.
-We further conducted ablation studies with results shown in Appendix.
-Quantitative Results
+To validate the value of Adv-BMT scenarios in downstream autonomous driving (AD) tasks, we train a reinforcement learning (RL) agent within augmented scenarios containing collision interactions generated by Adv-BMT and baseline methods.
+To determine the quality of the augmented training scenarios, we measure both the driving performance and the safety performance of the learned AD agent compared to learning on the original training set.
+In our experiment, we conduct two sets of training: (1) open-loop RL, where the agent is trained on a fixed training set with adversarial scenarios generated based on ground-truth ego trajectories; and (2) closed-loop RL, where an adaptive adversarial agent attacks the current ego agent based on its recent rollout trajectory records.
+The adaptive adversarial agent’s motion is generated by Adv-BMT or a baseline method.
+Results are shown in Table 4 and Table 5.
+Setting.
+The training set contains 500 real-world scenarios randomly selected from the WOMD training set.
+We train a Twin Delayed DDPG (TD3) agent for 1 million steps using 8 random seeds to ensure robustness in MetaDrive li2022metadrivecomposingdiversedriving (hyperparameters listed in the appendix).
+We measure the average reward, average step cost, average route completion rate (Compl.), and average episode cost (cost sum) for driving performance measurement.
 To evaluate the impact of adversarial training using Adv-BMT-generated scenarios, we assess policy performance across two distinct validation environments: (1) 100 Waymo validation environments, which consist of unmodified real-world driving scenarios from WOMD validation set, and (2) 100 Adv-BMT environments, which is the augmented collision scenarios from the 100 validation scenes.
-Table 3 presents the evaluation results, demonstrating how training with Adv-BMT improves policy resilience in both standard and adversarial settings.
-A detailed analysis of these results follows in the subsequent sections.
-Table 3: Open-loop learning evaluation.
+Table 4: Open-loop RL agent evaluation.
+Each WOMD scenario is augmented with one collision scenario, so that the ratio between real-world and safety-critical scenes is 1:1.
+For each method, we augment the original dataset into a new training set.
 (a) Evaluation in the Waymo validation environments
-Training Scenarios Reward $\uparrow$ Cost $\downarrow$ Compl. $\uparrow$ Coll. $\downarrow$ Cost Sum $\downarrow$
+Training Scenarios Reward $\uparrow$ Cost $\downarrow$ Completion $\uparrow$ Collision $\downarrow$ Cost Sum $\downarrow$
 Waymo ettinger2021largescaleinteractivemotion 32.03 $\pm$ 4.27 0.39 $\pm$ 0.07 0.72 $\pm$ 0.05 0.14 $\pm$ 0.02 1.41 $\pm$ 0.35
@@
 Adv-BMT 31.47 $\pm$ 3.21 0.38 $\pm$ 0.03 0.73 $\pm$ 0.04 0.11 $\pm$ 0.02 1.35 $\pm$ 0.40
-Adv-BMT (Bi) 33.22 $\pm$ 1.83 0.36 $\pm$ 0.03 0.74 $\pm$ 0.03 0.12 $\pm$ 0.02 1.39 $\pm$ 0.22
-(b) Evaluation in the Adv-BMT Environments
-Training Scenarios Reward $\uparrow$ Cost $\downarrow$ Compl. $\uparrow$ Coll. $\downarrow$ Cost Sum $\downarrow$
+Adv-BMT (Refined) 33.22 $\pm$ 1.83 0.36 $\pm$ 0.03 0.74 $\pm$ 0.03 0.12 $\pm$ 0.02 1.39 $\pm$ 0.22
+(b) Evaluation in the Adv-BMT validation environments
+Training Scenario Reward $\uparrow$ Cost $\downarrow$ Completion $\uparrow$ Collision $\downarrow$ Cost Sum $\downarrow$
 Waymo ettinger2021largescaleinteractivemotion 37.01 $\pm$ 6.16 0.64 $\pm$ 0.09 0.60 $\pm$ 0.07 0.30 $\pm$ 0.02 2.96 $\pm$ 0.63
@@
 Adv-BMT 37.33 $\pm$ 3.57 0.62 $\pm$ 0.03 0.63 $\pm$ 0.04 0.25 $\pm$ 0.05 2.41 $\pm$ 0.43
-Adv-BMT (Bi) 39.55 $\pm$ 2.94 0.59 $\pm$ 0.04 0.65 $\pm$ 0.02 0.27 $\pm$ 0.04 2.74 $\pm$ 0.54
-Table 4: Closed-loop learning results
+Adv-BMT (Refined) 39.55 $\pm$ 2.94 0.59 $\pm$ 0.04 0.65 $\pm$ 0.02 0.27 $\pm$ 0.04 2.74 $\pm$ 0.54
+Table 5: Closed-loop RL agent evaluation.
+We use the same augmented training set as in the open-loop experiment.
+For the adaptive adversarial learning experiment, we implement an adaptive generator for Adv-BMT and CAT zhang2023catclosedloopadversarialtraining .
+We discard the other baseline methods due to their low generation speeds.
 (a) Waymo Validation Environments
-Adaptive Generator Reward $\uparrow$ Cost $\downarrow$ Comp. $\uparrow$ Coll. $\downarrow$ Cost Sum $\downarrow$
+Generator Reward $\uparrow$ Cost $\downarrow$ Completion $\uparrow$ Collision $\downarrow$ Cost Sum $\downarrow$
 CAT 32.15 $\pm$ 2.89 0.38 $\pm$ 0.03 0.74 $\pm$ 0.04 0.10 $\pm$ 0.00 2.02 $\pm$ 0.24
@@
 (b) Adv-BMT Environments
-Adaptive Generator Reward $\uparrow$ Cost $\downarrow$ Comp. $\uparrow$ Coll. $\downarrow$ Cost Sum $\downarrow$
+Generator Reward $\uparrow$ Cost $\downarrow$ Completion $\uparrow$ Collision $\downarrow$ Cost Sum $\downarrow$
 CAT 39.47 $\pm$ 3.88 0.62 $\pm$ 0.05 0.63 $\pm$ 0.04 0.22 $\pm$ 0.02 2.51 $\pm$ 0.45
 Adv-BMT 40.40 $\pm$ 6.39 0.57 $\pm$ 0.04 0.63 $\pm$ 0.05 0.22 $\pm$ 0.04 2.48 $\pm$ 0.97
-Analysis
-1) Improved Policy Robustness in Adversarial Settings.
-Table 3 presents the evaluation results comparing policies trained with Adv-BMT-generated scenarios against baselines in both real-world and safety-critical validation environments.
-Adv-BMT demonstrates superior performance in reducing collision rates while maintaining high completion rates.
-In the collision validation environments, Adv-BMT-trained policies achieve the lowest collision rate while preserving competitive completion rates.
-This highlights the effectiveness of Adv-BMT in exposing policies to diverse and high-risk scenarios, leading to improved safety performance.
-Furthermore, Adv-BMT achieves the lowest sum cost, suggesting that policies trained on Adv-BMT-generated data are more efficient in handling safety-critical interactions.
-Besides, the results from adaptive generators shows an slight improvement in performance compared to training in fixed scenario sets.
-In the open-loop learning, among the baselines, STRIVE achieves the highest reward (37.72) in collision environments, but this comes at the cost of a higher collision rate (0.29).
-SEAL and CAT exhibit higher costs and collision rates, which suggests that adversarial training leads to suboptimal behaviors.
-In contrast, agent training with Adv-BMT attacks results in improvement on both safety and task completion.
-We conduct an ablation study on Adv-BMT plus the forward refinement method, and results indicate a slightly better performance in average rewards, and average cost compared to Adv-BMT.
-From closed-loop learning shown in table 4, we can see that agent learned with Adv-BMT environments has a significantly lower driving trajectory costs and higher rewards compared to baseline methods.
-With last step costs, collision rates, and driving completion rate remain consistent with the baseline.
-Comparing across tables with open-loop evaluations, we can see improvements over all metrics including rewards, completion rates, collision rates, and trajectory costs.
-This validates the value of diverse safety-critical scenarios which is provided in the closed-loop setting.
-5 Conclusion and Limitations
-Conclusion
-This work introduces Adv-BMT, a novel safety-critical scenario generation framework that leverages BMT model to bring diverse and realistic adversarial collisions into real-world traffic scenarios.
-Adv-BMT consists of three key components: (1) collision initialization, which samples a potential collision outcome as the last frame; (2) reverse prediction, which generates the adversarial pre-collision trajectories; (3) rule-based filtering, which ensures adversarial agents’ motions remain realistic and feasible.
-We evaluate Adv-BMT across multiple benchmarks and experiments and demonstrate that training with generated adversarial scenarios improves AD agent performance in both held-out log-replay and collision test sets with a clear margin.
-Limitations
-Following limiations of our current work will be potentially addressed in the future: (1) we only consider adversarial behaviors among vehicle agents, but not among other agent types like pedestrains and bicycles.
-(2) we only evaluate our adversarial scenarios in RL policies, but adversarial scenarios can also be leveraged in human-in-the-loop learning.
+Analysis.
+Table 4 reports open-loop evaluations comparing RL policies trained on Adv-BMT–generated scenarios against policies trained on baseline scenario sets.
+On the WOMD original validation environments, Adv-BMT–trained agents outperform all baselines across metrics: episode cost decreases by 10% and collision rate by 8%, while reward and route-completion remain comparable to the best baseline.
+Adding the optional forward refinement (Sec.
+3.3) yields additional gains—reward increases by 6%, episode cost decreases by 5%, and route completion improves slightly.
+On safety-critical test environments, Adv-BMT further reduces episode cost by 17% and collision rate by 14% relative to the best baseline.
+With forward refinement, we observe further improvements in driving capability, reflected in higher reward and completion rates.
+These results indicate that training on Adv-BMT scenarios exposes agents to a broader distribution of high-risk interactions, enabling more robust, safety-aware policies.
+The extra gains from forward refinement underscore the importance of modeling reactive traffic around emerging incidents, which better prepares agents for cascading hazards.
+Motivated by these findings, we next evaluate all methods in a closed-loop RL setting.
+Table 5 reports closed-loop RL results under reactive adversarial traffic generators.
+Relative to the corresponding open-loop policy, the closed-loop policy improves across all metrics in both real-world and safety-critical collision environments, indicating a substantially safer and more robust AD policy.
+Compared with the strongest closed-loop baseline (CAT), our method achieves higher reward, lower collision rate, and a 38% reduction in episode cost.
+These gains suggest that training with Adv-BMT scenarios—featuring diverse adversarial behaviors and realistic, traffic-consistent reactions—better train the agent to anticipate and mitigate high-risk interactions.
+Learning in closed-loop further improves performance; by training the agent to respond online to reactive opponents, it becomes better at anticipating dangerous situations, and staying safe in unfamiliar or rapidly changing traffic conditions.
+The results also highlight the importance of the Adv-BMT design, which enables the generation of realistic and flexible multi-agent interactions with the adversarial agent.
+Figure 6: Adv-BMT can also generate scenarios of pedestrian and cyclist adversarial agents.
+Figure 7: Baseline method comparisons.
+We present several examples where Adv-BMT successfully generates collision interactions in scenarios where all baseline methods fail.
+Unlike the baselines, Adv-BMT does not rely on the proximity of neighboring agents to the ego vehicle, enabling more flexible and diverse adversarial attack strategies.
+5 Conclusion
+We introduced Adv-BMT, a novel framework for generating diverse safety-critical driving scenarios with realistic traffic interactions.
+Built on a Bidirectional Motion Transformer with the ability to perform bidirectional motion prediction tasks, Adv-BMT is able to design a new adversarial agent for each real-world scenario.
+A key advantage of Adv-BMT lies in its ability to generate realistic and flexible multi-agent interactions surrounding the adversarial agent.
+Unlike previous frameworks that select an existing neighboring vehicle and modify the corresponding trajectory for an adversarial attack, Adv-BMT initializes candidate collisions for a new adversarial agent, reconstructs multi-agent trajectories via inverse prediction, and optionally applies forward refinement for reactive, traffic-consistent interactions.
+Adv-BMT is able to balance realism, controllability, and scene reactivity.
+Our evaluations validate the quality of Adv-BMT scenarios in terms of interaction realism and diversity; furthermore, we demonstrate that learning within the Adv-BMT traffic flow improves AD agents’ safety performance by a clear margin.
+Limitations.
+Adv-BMT relies on long token sequences to represent multi-agent traffic scenes, which results in high memory and computational demands.
+Our empirical study focuses on a single downstream task—reinforcement learning for an autonomous driving agent—so broader applicability remains to be validated.
+Acknowledgment: This work was supported by the NSF Grants CNS-2235012, IIS-2339769, and CCF-2344955. ZP was supported by the Amazon Fellowship via UCLA Science Hub.
 Appendix A Model Details
 A.1 BMT Architecture
-Scenario Format and Preprocessing
-The model takes real-world scenarios with the format from ScenarioNet li2023scenarionetopensourceplatformlargescale , which unifies real-world datasets including caesar2020nuscenesmultimodaldatasetautonomous  caesar2022nuplanclosedloopmlbasedplanning  wilson2023argoverse2generationdatasets .
-The model performs a series of preprocessing on both static map features and dynamic agent trajectories.
-We compute the global boundary to extract a consistent map center and heading.
-Each map feature is decomposed into a sequence of vectorized segments, where each vector is represented by a start and end point in 3D coordinates.
-These vectors are then augmented with directional relative positions, segment headings, and lengths.
+Preprocessing.
+The model takes real-world scenarios in the ScenarioNet format li2023scenarionetopensourceplatformlargescale .
+It performs a series of preprocessing steps on both static map features and dynamic agent trajectories.
+We compute the global bounds to extract a consistent map center and heading.
+Each map feature is decomposed into a sequence of vectorized segments, where each vector is represented by start and end points in 3D coordinates.
+These vectors are augmented with relative directional positions, segment headings, and lengths.
 Semantic attributes are binary indicators that encode feature types.
-The map feature types (lane, crosswalk, broken line, yellow line, stop sign, etc.) are encoded for semantic information.
-Features are then centralized to map center.
-Traffic light states are also encoded and aligned with the steps.
-Similarly, agent trajectories are also centralized for agent feature encode, with information of (position, heading, velocity, shape, and type) extracted across time for temporal sequences.
-Besides, there is reordering for ego agent and agents to be predicted.
+Map feature types (e.g., lanes, crosswalks, broken lines, yellow lines, stop signs, etc.) are encoded as semantic indicators.
+Features are then centered at the map center.
+Traffic-light states are also encoded and aligned with timesteps.
+Similarly, agent trajectories are centered for agent-feature encoding, and position, heading, velocity, shape, and type are extracted over time to form temporal sequences.
+We also reorder the ego agent and target agents to fixed indices.
 We extract a 16-dimensional state vector at each timestep.
-The preprocessing ensures all trajectories are represented in a consistent spatial frame and translated to an ego-centric coordinate system.
-Tokenization
-We formulate motion prediction as an auto-regressive next-token prediction problem, where each motion token corresponds to a discretized control input for a fixed time interval.
-Tokenization process maps continuous agent motion (acceleration and yaw rate) into discrete 2D bins.
-The tokenization process considers candidate tokens sampled from a fixed grid of bins, and simulates the resulting motion over a short duration, and then selects the best-matching action by minimizing the contour alignment error between the predicted agent shape and the ground-truth position and heading.
-We adopt a simplified version of the bicycle model to parameterize agent motion using longitudinal acceleration and yaw rate within predefined bounds: acceleration is limited to the range of $[-10,10]$ m/s2, and the yaw rate is constrained to $[-\pi/2,\pi/2]$ rad/s.
-With predicted motion token sequences, the trajectory can be reconstructed by mapping the token back to the acceleration and yaw angle change.
-In both forward and reverse directions, motion tokens are decoded into continuous trajectories using the same tokenizer.
-Given an initial state $\tau_{t}=(x_{t},y_{t},\theta_{t},v_{t})$, the forward decoding process simulates the agent’s next state $\tau_{t+1}$ by applying a tokenized control action $z_{t}=(a_{t},\delta_{t})\in\mathcal{A}$, where $\mathcal{A}$ denotes the discrete token space.
-This is repeated autoregressively over the sequence of predicted tokens to reconstruct the full trajectory $\boldsymbol{\tau}_{t:t+T}$.
-In contrast, reverse decoding starts from a known future state $\tau_{t+1}$, the model evaluates all possible token candidates $z_{t}\in\mathcal{A}$, simulates their inverse dynamics using $\Delta t\to-\Delta t$, and selects the token that best reconstructs the preceding state $\tau_{t}$.
-The ability to operate in both directions is a key distinction of our approach: forward prediction enables open-loop simulation of future behaviors, while reverse prediction allows us to trace back from a desired outcome (e.g., a collision state) to plausible initiating actions.
-Decoding Architecture.
-The decoder follows a GPT-like structure composed of stacked cross-attention layers.
-Each layer integrates three structured attention modules: agent-to-agent (A2A), agent-to-temporal (A2T), and agent-to-scene (A2S).
-These modules attend over dynamically constructed graphs defined by spatial or temporal adjacency.
-Relational information across modalities is captured via multiple embeddings: Agent-to-Agent Relation Embedding, Agent-to-Time Relation Embedding, and Agent-to-Scene Relation Embedding, each encoding context-specific spatial information.
-For token construction, several embeddings are used, including the Agent Type Embedding, Agent Shape Embedding, Agent ID Embedding, and the Motion Token Embedding which maps discrete control tokens.
-A Continuous Motion Feature Embedding is applied to embed acceleration and yaw rate attributes.
-Auxiliary embeddings include the Special Token Embedding for indicating sequence boundaries, and the Backward Prediction Indicator Embedding to distinguish between forward and backward prediction modes.
-For each attention edge, a relative relation embedding is computed using a Fourier encoder and added to the key and/or value vectors.
-The input agent token $\mathbf{X}\in\mathbb{R}^{B\times T\times N\times d}$ (batch size $B$, time steps $T$, number of agents $N$, hidden dimension $d$) is progressively updated across layers by aggregating contextual features from other agents, their temporal history, and relevant map elements.
-Motion Token Embedding
+This preprocessing ensures that all trajectories are represented in a consistent spatial frame and expressed in an egocentric coordinate system.
+BMT Motion Tokenization.
+We formulate motion prediction as an autoregressive next-token prediction problem, where each motion token corresponds to a discretized control input over a fixed time interval.
+The tokenization process maps continuous agent motion—longitudinal acceleration and yaw rate—into discrete 2D bins.
+During tokenization, we consider candidate tokens sampled from a fixed grid; for each candidate, we simulate the resulting motion over a short duration and select the best-matching action by minimizing the contour-alignment error between the predicted agent footprint and the ground-truth pose (position and heading).
+We adopt a simplified bicycle model to parameterize agent motion using longitudinal acceleration and yaw rate within predefined bounds: acceleration $\in[-10,10]\,\mathrm{m/s^{2}}$ and yaw rate $\in[-\pi/2,\pi/2]\,\mathrm{rad/s}$.
+Given predicted motion-token sequences, the trajectory is reconstructed by mapping tokens back to acceleration and yaw-rate changes.
+In both forward and reverse directions, the same tokenizer is used to decode motion tokens into continuous trajectories.
+Given an initial state $\tau_{t}=(x_{t},y_{t},\theta_{t},v_{t})$, forward decoding simulates the next state $\tau_{t+1}$ by applying a tokenized control action $z_{t}=(a_{t},\delta_{t})\in\mathcal{A}$, where $\mathcal{A}$ is the discrete token space; repeating this autoregressively over the predicted tokens reconstructs the full trajectory $\boldsymbol{\tau}_{t:t+T}$.
+For reverse decoding, given a known future state $\tau_{t+1}$, the model evaluates candidate tokens $z_{t}\in\mathcal{A}$, simulates their inverse dynamics with $\Delta t\to-\Delta t$, and selects the token that best reconstructs the preceding state $\tau_{t}$.
+Operating in both directions is a key distinction of our approach: forward prediction enables open-loop rollouts of future behaviors, whereas reverse prediction traces from a desired outcome (e.g., a collision state) back to plausible initiating actions.
+BMT Motion Decoder.
+The decoder follows a GPT-like structure composed of stacked cross-attention blocks.
+Each block integrates three structured attention modules:agent-to-agent (A2A), agent-to-temporal (A2T), and agent-to-scene (A2S).
+Relational information across modalities is encoded via relation embeddings: agent-to-agent relation embedding, agent-to-time relation embedding, and agent-to-scene relation embedding, each capturing context-specific spatial/temporal cues.
+For token construction, we use several embeddings, including the agent-type embedding, agent-shape embedding, agent-ID embedding, and motion-token embedding (which embeds discrete control tokens).
+A continuous-motion feature embedding encodes acceleration and yaw-rate attributes.
+Auxiliary embeddings include the special-token embedding (to indicate sequence boundaries) and the reverse-prediction indicator embedding (to distinguish forward from reverse prediction modes).
+For each attention edge, a relative-relation embedding is computed using a Fourier encoder tancik2020fourierfeaturesletnetworks and added to the key and/or value vectors.
+The input agent tokens are progressively updated across layers by aggregating contextual features from other agents, their temporal histories, and relevant map elements.
+A.2 BMT Motion Decoding
+Input.
 Specifically, each input agent token to the BMT motion decoder is constructed by summing of embeddings of:
@@
 as well as a continuous motion delta feature embedded via a Fourier encoder. These components are projected into the same hidden dimension and summed to form the input motion token embedding.
-Prediction Heads and Outputs
+Prediction Head and Output.
 After processing through all decoding layers, the final hidden state for each valid token is passed through a two-layer MLP head to produce logits over the motion token space:
@@
 where $\phi(\cdot)$ denotes the GELU activation and $K^{2}$ is the number of discrete motion tokens (from a $K\times K$ acceleration–yaw bin grid).
-The resulting logits are used to predict the next motion token at each time step. During inference, we generate motion tokens using nucleus (top-$p$) sampling.
+The resulting logits are used to predict the next motion token at each time step.
+During inference, we generate motion tokens using nucleus (top-$p$) sampling.
 Trajectory Reconstruction
-Our model makes a prediction in the interval of 5 time steps (0.5 seconds).
+Our model makes a prediction in the interval of 0.5 seconds.
 To simulate the effect of a motion token over a fixed time step $\Delta t=0.5\,\mathrm{s}$, we adopt midpoint integration based on a simplified bicycle model.
@@
 $x_{t}=x_{t+1}-\bar{v}\cdot\cos(\bar{\theta})\cdot\Delta t,\quad y_{t}=y_{t+1}-\bar{v}\cdot\sin(\bar{\theta})\cdot\Delta t.$
-Decoder Training Loss
+A.3 BMT Training
+Training Loss.
 The decoder produces a logit tensor $\hat{\mathbf{z}}\in\mathbb{R}^{B\times T\times N\times|\mathcal{A}|}$, where $|\mathcal{A}|$ is the number of motion tokens (i.e., discretized acceleration–yaw pairs).
@@
 where CE denotes the standard cross-entropy loss between the predicted logits and the ground-truth discrete token.
-Reverse Prediction
+Reverse Prediction.
 During reverse prediction, the model measures metrics separately for forward and reverse token predictions.
@@
 with analogous expressions for forward prediction (i.e., for $1-b_{b,t,n}$).
-Metrics
+Metrics.
 To measure the quality and diversity of the model’s predictions during training, we track the perplexity:
@@
 $\text{Cluster}=\sum_{a\in\mathcal{A}}\mathbf{1}[\bar{p}_{a}>0].$
-Total Loss
+Total Loss.
 The total loss is the sum of all enabled components:
@@
 Total 5,287,105 20.17
-A.2 Training Details
+A.4 Training Details
 Our model has 5.2 million trainable parameters, with detailed break downs indicatd in Table 6.
@@
 Figure 8: Diverse adversarial behaviors generated by Adv-BMT.
-Appendix B Experiment Details
-Training Environment
+Appendix B Additional Experiment Results
+Training Environment.
 We conduct our reinforcement learning experiments using the MetaDrive ScenarioEnv li2022metadrivecomposingdiversedriving , which provides standardized driving environments for training and evaluating autonomous agents.
@@
 Driving reward is measured by forward lane progress, while penalties are applied for collisions with other vehicles or drifting off-road.
-Hyper-parameters
+Hyperparameter.
 The settings of our open-loop and closed-loop adversarial RL experiments are shown in table 8.
@@
 Appendix C Quantitative Results
-C.1 Waymo Open Sim Agents Challenge
+C.1 Ablation Study
+Table 9: Realism Results.
+Method $_{\text{avg}} SFDE$$_{\text{min}}$ $_{\text{avg}} SADE$$_{\text{min}}$ $_{\text{agent,min}} Coll$$_{\text{agent,avg}}$
+Reverse 3.92 2.69 5.53 5.11 0.03 0.05
+Reverse + Adv-init 7.17 5.56 7.68 7.12 0.13 0.16
+Reverse + Adv-init + Filter 6.93 5.30 7.58 7.02 0.12 0.15
+Table 10: Diversity Results.
+[…]
+Reverse 8.35 3.13 0.17 0.71 0.02
+Reverse + Adv-init 9.89 3.82 0.22 0.64 0.13
+Reverse + Adv-init + Filter 10.27 3.99 0.22 0.73 0.09
+Reverse predictions with adversarial initializations exhibit greater deviation from the ground-truth data and yield improved diversity.
+This behavior is expected, since adversarial initializations modify the terminal positions and headings of selected agents, which propagates backward into more varied histories.
+Importantly, our rule-based filtering does not reduce diversity; rather, it preserves multimodality while improving realism metrics by removing physically implausible trajectories.
+C.2 Waymo Open Sim Agents Challenge
 Table 11: WOSAC Evaluation results of BMT.
@@
 Metametric 0.554 0.753
-We evaluated BMT resutls on 400 WOMD validation scenarios using Waymo Open Sim Agents Challenge (WOSAC) 2024 montali2023waymoopensimagents .
+We evaluate BMT results on 400 WOMD validation scenarios using the Waymo Open Sim Agents Challenge (WOSAC) 2025 montali2023waymoopensimagents .
 Evaluation results are summarized in Table 11.
-Within the metrics, for minADE metrics, smaller values indicate better more accurate perdictions, and for the rest metrics, the larger score means better performance.
-Within the results, we can see that forward prediction achieves much better performance compared to reverse prediction across all metrics, except similar performance in Angular speed, angular acceleration, distance to nearest object, and TTC.
-In these mentioned metrics.
-Note that the training time for forward prediction and reverse prediction are similar (10E6 and 15E6 steps respectively).
-The WOSAC results indicate that our BMT model is better at predicting future motion than predicting history motion.
-We will submit to the official WOSAC learderboard for further reference in the future.
+For the metrics, smaller values of minADE indicate more accurate predictions, whereas larger values for the remaining metrics indicate better performance.
+From the results, we observe that forward prediction achieves much better performance than reverse prediction across all metrics, except for similar performance in angular speed, angular acceleration, distance to the nearest object, and TTC.
+Note that the training times for forward prediction and reverse prediction are similar (10E6 and 15E6 steps, respectively).
+The WOSAC results indicate that our BMT model is better at predicting future motion than historical motion.
+Upon reviewing the collisions detected in reverse predictions with real initializations, we found that most occurred in crowded parking-lot areas, where clusters of parked vehicles or pedestrians are close together.
+These were flagged as collisions by WOSAC’s evaluation API, even though they may not represent meaningful agent–agent interactions.
+This explains why metrics such as ADE and FDE remain comparable across methods despite differences in collision scores.
+The performance gap is primarily due to our training strategy: the model was pre-trained for forward prediction (around 800k steps) and then fine-tuned for bidirectional prediction (around 1.5M steps).
+We have included these details in Section A.4 (Training Details).
 Appendix D Qualitative Results
-Visualizations
-In Fig.
-6, we demonstrate six pairs of qualitative results by Adv-BMT.
-In different scenarios, our opponent behavior exihibits in different category of safety-critical driving patterns.
-This means that our augmented opponent agents have plausible interactions with realistic driving trajectories.
-Besides, in Fig.
-7, we demonstrate diverse collisions from just one single scenario.
-This is the power of Adv-BMT compares to other baseline work, which generates the same or similar adversarial behavior with the same driving log.
-Figure 7: Diverse adversarial behaviors generated on one driving log by Adv-BMT.
-Demo Video
+Visualizations.
+Figure 9 presents six pairs of qualitative results generated by Adv-BMT.
+Across different scenarios, the adversarial agents exhibit a diverse range of safety-critical driving behaviors, demonstrating their ability to interact plausibly with realistic traffic participants.
+The visualizations illustrate that Adv-BMT can generate multiple distinct collision outcomes from a single driving log.
+This highlights a key advantage of Adv-BMT over baseline methods, which tend to produce identical or highly similar adversarial behaviors for the same input scenario.
+Demo Video.
 We submit a video within our supplementary materials.
-Here we put visualizations with case studies through animated visualizations of Adv-BMT scenarios, which simulates different types of vehicles, pedestrians and bicycle agents.
+Here we provide visualizations with case studies through animated simulations of Adv-BMT scenarios, which include different types of vehicles, pedestrians, and bicycle agents.
+More demos are available at https://metadriverse.github.io/adv-bmt/.
+Figure 9: Diverse adversarial behaviors generated by Adv-BMT.
 Appendix E Broader Impacts